A Complete Bibliography of Proceedings of the ACM on Measurement and Analysis of Computing Systems

(POMACS)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

26 May 2022 Version 1.03

Title word cross-reference

1/N [Gas17]. 3 [KKRK22, LKJK20, LGC^{+}18]. $9.81 [FS20]. D [YXL21, HV21]. k [CKMP19, SGHB20].

-Accurate [Gas17]. -Core [CKMP19]. -hops [YXL21].

1 [CP18, SHBSW20].

2021 [CCR21]. 2022 [CCR22]. 256 [BRC21].

5G [PLX22]. 5G-LTE [PLX22].

Accelerating [Cas17]. Access [LYE19, VPK18, WM20, WXW22]. Access-Delay [LYE19]. Accesses [DMB^{+}22]. Accurate [BNS17, Gas17, JHWC21]. Achievable [RB20]. Achieving [WW20, YCX^{+}19]. Acoustic [IYRR18]. Across [SATN22]. Ad [SVL20, TMM20]. Adaptation [WCX^{+}18]. Adaptive [SBY19]. Additional [YPA^{+}18]. additive [WS17a]. Addresses [MLHC20]. adPerf [PBLC21]. Ads [PBLC21]. Adversarial [CSX17, LSNI21, LM18]. Adversary [RKTV22]. Adversary-resistant [RKTV22]. Advertising [PPL^{+}21, YWB19]. Age [SHBSW18]. Age-Based [SHBSW18].

Agent [LVW19, SGS19]. Aggregate [PTD20]. Aggregation [KLCS19, LCS^{+}21]. Agnostic [AASS18]. Agreements [WM20]. Ahead [SLJ19]. Algorithm [DCKG19, JIAM19, TTL^{+}21, YWH17, YDH^{+}18]. Algorithmic [NT19, VS19]. Algorithms [CMM22, CYC^{+}19, CCG21, FKP^{+}21, HCZ^{+}21, KMM21, LNY21, SZL^{+}20, YZH^{+}21, ZWT^{+}17]. Alignment [CKMP19, DCKG19]. Allocation [RKTV22, TSLG^{+}20, ZLW17]. Allocations [TP18]. Alternative [FS19]. Among [PJDQ20]. Analysis [AASS18, ALR17, BRC21, BNS17, CYG^{+}17, CMH^{+}17, DCKG19, HBV19, HSV19, LKS^{+}17, LML^{+}20, NVGT20, OKB^{+}18, PSC19, SAM^{+}22, SHBSW18, SSM^{+}17, WM20, YHG^{+}17]. Analytics [AEKB20, KLCS19]. Analyzing [YWB19]. Android [ZYH^{+}19]. Anonymity [FVB^{+}18, VFV17]. App [ZYH^{+}19]. Application [SZL^{+}20, YNJS21, ZYH^{+}19]. Applications [BZY^{+}22]. Approach [AdV18, AGS17, BCM17, FZCL18, GTL^{+}17, KLCS19, NT19, SSM^{+}17, TSLG^{+}20, WKWY19, ZZLL21]. Approximate [KKT^{+}19, PHSB21, WLS19]. Approximation [DE20, Gas17, GV17, HCZ^{+}21, Zho22]. Approximations [AG22, RBT21, Yin17]. Apps [ZTY^{+}21]. Architecture [KKT^{+}19, LKJK20, LYJ^{+}21]. Architecture-Aware [KKT^{+}19]. Architectures [GFL^{+}22]. Argus [RS22]. Armed [SGS19, YPA^{+}18]. Arms [CPF20]. Assessment [NT19]. Assisted [GKK20]. Asymptotic [Cas17, CMM22, DV18, KV21, WW20]. Asymptotically [MBvL18]. Attacker [WS17a]. Attacks [FM19, IYRR18]. Attraction [Van19]. Auctions [KMM21, TMM20]. Augmentation [RS22]. Automatic [SMF^{+}22]. Autonomous [ZHZ^{+}21]. Aware [BSA^{+}21, KKT^{+}19, KKRK22, LYJ^{+}21, VPK18]. Axiomatizing [ZMSS19].

Backbone [SAM^{+}22]. Backhaul

[YHG^{+}17]. Balanced

[BCM17, LGH^{+}17, PJDQ20, YLX17]. Balancing [ALR17, GSHB19, HBV19, HV21, HSV19, MCS^{+}19, MBvL18, TS19, WZS20, ZWT^{+}17, ZWT^{+}18, ZTS18]. Bandit [CES19, YPA^{+}18, ZZLL21]. Bandits [CPF20, CCG21, DV18, SGS19]. Bankrupting [PPL^{+}21]. Based [AdV18, NGB^{+}21, SHBSW18, TS19, YWB19, ZZCC17, CMH^{+}17, CAJK17, FM19, KLCS19, SAM^{+}22, Van19, ZTS18]. Batch [AI19]. Behavior [CMM22, HWW^{+}20]. Behind [ASV^{+}21]. Below [HSV19]. Benchmark [HHP20, WZT19]. Best [SZL19]. Better [CJC^{+}19, GTL^{+}17]. between [WM20]. BGP [SMF^{+}22]. Bidding [KMM21, TMM20]. Big [ZZCC17]. Billion [KCT18, NNVD17]. Billion-Scale [KCT18, NNVD17]. Bitcoin [VFV17]. Blockchain [GSWV20, HWW^{+}20]. Blockchains [DAR^{+}21, RKTV22]. Blocking [SVL20]. Bonus [SCYO21]. Bootstrapped [EC18]. Bottleneck [RGBY^{+}19, ZMW^{+}20]. Bought [FS20]. Bound [CP18]. Bounded [KCT18, SZL19]. Bounds [BBHB18]. Bregman [AN20]. Bregman-style [AN20]. Browsers [PSC19]. Buffer [CPR19, YWH17]. Burstable [WUNK17]. Byzantine [CSX17].

Cache

[QTES19, WXW22, ZKAV20, AMA18].

Caching [AMA18, AGS17, BBHB18, BBS20, LSNI21, LYE19, TQJS18]. Cake [KB20]. Calculate [JIAM19]. Calculus [BNS17]. Canonical [DCKG19]. Capabilities [WZT19]. Carlo [LKE19]. Cascade [NNVD17]. Cascades [HC19]. Case [FWB17, GKK20, KKRK22]. Catchment [SK19]. Causal [PSC19, ZL18]. Cause [LML^{+}20]. Cellular [XJ22, YHG^{+}17]. Censored [AS17]. Censorship [JHWC21]. Census [MSJ^{+}19]. Centaur [LKJK20]. Centers [AAJV20, IYRR18, ZCW21]. Centric [PJDQ20]. Cerberus [GZB^{+}21]. Chain [CT19]. Channel [IYRR18, TWFO20]. Characterization [BZY^{+}22, CYG^{+}17, CMM22, LKS^{+}17]. Characterizing [CLS^{+}21, PBLC21, SvKB^{+}20, WLdA^{+}17, XWG^{+}21, ZMW^{+}20]. Charging [SZL^{+}20]. Chasm [ZHM^{+}21]. China [ZMW^{+}20]. Chipotle [FS20]. Chips [LKS^{+}17]. Choices [GZB^{+}21, HV18]. Circuit [ZCW21]. Class [CAJK17, JNTG18]. Classification [ASV^{+}21]. Clean [SHBSW18]. Clocks [TL20]. Close [TLL^{+}18]. Closed [Cas17]. Cloud [CYC^{+}19, IKS21, PJDQ20, PG17, SIS17, WUNK17, ZLW17]. Cloud-Centric [PJDQ20]. Codes [MCS^{+}19]. Coding [WLS19]. Collaboratively [SZL19]. Collection [ASME19]. Combinatorial [CCG21]. Communication [DBS17, FKP^{+}21, LVW19]. Communities [SMF^{+}22]. Compact [SQ20]. Comparative [PLX22]. Competitive [LYP^{+}19, SLJ19, SZL^{+}20, YZH^{+}21]. Compiler [SDD^{+}17]. Compiler-Inserted [SDD^{+}17]. Compilers [XLY^{+}22]. Complex [GLH^{+}19]. Complexity [AGGS20, AGS17, WS17b, ZWT^{+}17]. Comprehensive [SATN22]. Computation [DAR^{+}21]. Computationally [CT19]. Compute [JJS17]. Computers [SIR20]. Computing [KKT^{+}19, LGH^{+}17, TKZ^{+}18]. Concealing [SLWS20]. Conditions [SGHB20, ZTS18]. Configuration [KCMX20]. Configurations [SDA18]. Congested [WMX17]. Congestion [MSJ^{+}19, RGBY^{+}19, ZMSS19]. Congestion-Controlled [RGBY^{+}19]. Connected [ALBS^{+}20]. Connectivity [SAM^{+}22]. Consequences [VS19].

Consistency [LYQ^{+}22]. Constant

[CMM22]. Constrained [KMM21, LCS^{+}21]. Constraint [HCZ^{+}21, TTL^{+}21]. Constraints [AN20, AGS17, LYP^{+}19, WYN20, WZS20, YHS^{+}20]. Consumption [BRC21, CNW^{+}17]. Contemporary [JKR18]. Content [GKK20, WM20]. Continuous [CPF20]. Contract [DAR^{+}21]. Contracts [CLS^{+}21, KMM21]. Contractual [WM20]. Control [AMSS19, AEKB20, CES19, DV18, LCS^{+}21, LYQ^{+}22, LGK22, MSJ^{+}19, WLdA^{+}17, ZMSS19]. Controlled [RGBY^{+}19]. Convection [KKRK22]. Convergence [DBS17, WS17b, WYN18]. Convex [AN20, SLJ19, YDH^{+}18, LGW20]. Cooperative [AGS17]. Core [CKMP19]. Correlated [DCKG19]. Cost [AMA18, AI19, BNS17, BSA^{+}21, HC18, PSLW22]. Cost-Aware [BSA^{+}21]. Cost-Efficient [AMA18]. Counterfeit [GWX^{+}20]. Coupled [WYN18]. CPU [RVC^{+}19]. Criteria [TMM20]. Critically [HV21]. Cross [OKB^{+}18]. Cross-layer [OKB^{+}18]. Crossbar [GTL^{+}17]. Crowdsourced [SVL20]. Crowdsourcing [SCYO21]. Cryptocurrency [FVB^{+}18, GWX^{+}20, TWFO20]. Crystal [ASME19]. CSI [SDD^{+}17]. Curing [HC18]. Curtain [ASV^{+}21]. Curvature [SAM^{+}22]. Curvature-based [SAM^{+}22].

D [LKJK20, LGC^{+}18, HV18, KKRK22]. Dandelion [FVB^{+}18, VFV17]. Data [AAJV20, AEKB20, ATN^{+}22, BSA^{+}21, CEF^{+}20, CAJK17, FS19, GGYZ17, HKY^{+}19, HCZ^{+}21, IYRR18, KKRK22, KLCS19, SIR20, TKZ^{+}18, TPK^{+}18, TKK21, ZKAV20, ZCW21, ZZCC17]. Data-Driven [ATN^{+}22]. Datacenter [GZB^{+}21, ZLM^{+}18, ZL18]. Datacenter-scale [ZL18]. DBMSes [SATN22]. DCN [WCX^{+}18]. DCNs [YCX^{+}19]. Debloating [KCMX20].

DECAF [IKS21]. December [CCR21]. Decentralized [XWG^{+}21]. Decision [WYN18]. Deconstructing [CNW^{+}17]. Deep [WCX^{+}18, XLY^{+}22]. Degradation [NT19]. Degree [ZWT^{+}18]. Delay [BZ18, LYE19, PSLW22, WS17b, WW20, ZWT^{+}17, ZWT^{+}18, ZTS18, Zub20]. Delay-Optimal [ZWT^{+}17, Zub20]. Delays [DBS17]. Demand [AS17, SSB18]. Demystify [ZYH^{+}19]. Demystifying [GLH^{+}19]. Density [LKJK20]. Dependent [HBV19, ZHZ^{+}21]. Deployment [BSA^{+}19]. depth [CMH^{+}17]. Descent [CSX17, SX19, WYN20]. Design [BNS17, GZB^{+}21, LKS^{+}17, TSLG^{+}20, ZCW21]. Design-Induced [LKS^{+}17]. Designing [CJC^{+}19, YHG^{+}17, ZWT^{+}17]. Desktop [SIR20]. Detailed [GYG^{+}18]. Detecting [CLS^{+}21, SLWS20, XWG^{+}21]. Detection [WKWY19]. Deterministic [BNS17]. Device [LYJ^{+}21]. Devices [BRC21, CYG^{+}17]. Diagnosis [ZL18]. Did [IBL^{+}19, BRC21]. Differentially [Zho22]. Diffusing [YLX17]. Diffusions [EC18]. Dimensional [LML^{+}20, SX19, AMSS19]. Direct [WXW22]. Directed [LKE19]. Discretization [SBY19]. Disparities [ZHM^{+}21]. Disruptions [IMMP20]. Dissecting [IKS21]. Distributed [AGS17, CSX17, DBS17, FKP^{+}21, HNS19, JIAM19, LVW19, MCS^{+}19, MDBvL17, RKTV22, SX19, SDA18, SSM^{+}17, KLCS19]. Distribution [NVGT20]. Distributions [CPR19]. Diverse [MPIZ17]. Does [NKY^{+}18]. Drain [CMH^{+}17]. DRAM [CYG^{+}17, GKK20, GYG^{+}18, GLH^{+}19, LKS^{+}17]. DRAMs [KKRK22]. Driven [ASME19, ATN^{+}22, KKRK22, SIS17, WCX^{+}18]. Driving [ZHZ^{+}21]. DSM [GKK20]. Dual [WYN20]. during [LYL^{+}22]. Dynamic [FKP^{+}21, FZCL18, IMMP20, KCT18, LGK22, RBT21, SQ20, TPK^{+}18]. Dynamical [LGK22]. Dynamics [FWB17, PPL^{+}21].

Early [LGC^{+}18]. Eat [KB20]. ECI [AMA18]. ECI-Cache [AMA18]. Edge [AEKB20, LGH^{+}17]. Editorial [BD19, CCR21, CCR22, CGZ17, GKW21, TM20]. Editors [WA18]. Effective [LGH^{+}17, YWB19]. Efficiency [SVL20]. Efficient [AMA18, CT19, CCG21, GFL^{+}22, HNS19, YCX^{+}19]. Elasticity [MDBvL17]. Electric [SZL^{+}20]. Embedding [HNS19]. Empirical [SATN22]. Enabling [CMH^{+}17]. Encrypted [ASV^{+}21, BSA^{+}19]. End [BZY^{+}22, GWX^{+}20, JHWC21, WXW22]. End-to-End [JHWC21, BZY^{+}22, GWX^{+}20]. Endurance [AMA18]. Energy [AN20, CNW^{+}17, CMH^{+}17, RVC^{+}19]. Energy-Harvesting [AN20]. Enhancing [ALBS^{+}20, CAJK17, TKK21]. Enough [LYJ^{+}21]. Entanglement [NVGT20]. Environment [LML^{+}20]. Environments [NKY^{+}18]. EOSIO [HWW^{+}20]. Epidemic [HC19]. Epidemics [HC18]. Episodic [SBY19]. Equilibrium [WM20]. Erdos [CKMP19, DCKG19]. Estimated [Gas17]. Estimation [AASS18, AS17, CT19, NNVD17, NKY^{+}18, YNJS21]. Ethereum [CLS^{+}21]. Evaluation [BNS17, JKR18]. Event [CMH^{+}17]. Event-based [CMH^{+}17]. Everything [NBB^{+}19]. Exchange [XWG^{+}21]. Executing [AI19]. Exfiltration [SIR20]. Expected [Gas17]. Experience [BSA^{+}19, GGYZ17]. Experimental [CYG^{+}17, GYG^{+}18, GLH^{+}19]. Explained [OKB^{+}18]. Explaining [ZHM^{+}21]. Exploiting [CAJK17, SIR20]. Exposing [EC18]. Extensions [CP18]. Extra [YLX17]. Extreme [PLX22]. Extremely [SGHB20].

Failure [CJC^{+}19, ZLM^{+}18]. Fair

[JIAM19, TP18]. Fairness

[BCM17, IMMP20]. Families [ZZM^{+}19].

Fast [BNS17, LML^{+}20, SQ20, WS17b]. Fast-Convergence [WS17b]. Fault [SDA18]. Fault-Tolerant [SDA18]. FCFS [GYSHB21]. Federated [ZZLL21]. Feedback [CES19, PSLW22, vdBBvL19]. Fi [VPK18]. Fiedler [DE20]. Field [AG22, Gas17, GV17, HSV19, SSM^{+}17, Van19, Yin17, RBT21]. Filters [SVL20]. Finding [MLHC20]. Fine [CMH^{+}17, XJ22]. Fine-Grained [XJ22, CMH^{+}17]. Fingerprints [ALBS^{+}20]. Finite [CPR19, DBS17]. Finite-Buffer [CPR19]. Finite-Sum [DBS17]. First [KMM21, PLX22]. Fit [NKY^{+}18]. Flash [CAJK17, LGC^{+}18]. Flash-based [CAJK17]. Flexible [QTES19]. Flow [JIAM19, QTES19]. Forecasting [FS19]. Fork [Zub20]. Fork-Join [Zub20]. Formal [FVB^{+}18]. Fragmentation [SSB18]. Framework [DMB^{+}22, SDD^{+}17, ZHZ^{+}21]. Free2Shard [RKTV22]. Function [Zho22]. Fundamental [BBS20, FM19, WLS19]. Fusing [LYL^{+}22]. Fuzzing [HHP20].

G [CP18, SGHB20, SHBSW20]. Game [BZY^{+}22, FZCL18, WS17a]. Game-Theoretic [FZCL18]. Gaming [IKS21]. Gap [CT19]. Garbage [ASME19]. Gazer [ASME19]. GB [BRC21]. General [SGHB20]. Generalized [ZZM^{+}19]. Geo [KLCS19]. Geo-distributed [KLCS19]. Geographic [AGS17]. GI [CP18]. GI/G/1 [CP18]. Gittins [SGHB20]. Global [JHWC21, Van19]. go [BRC21]. Gossiping [ZZLL21]. GPGPU [YNJS21]. GPU [JKR18, KKRK22, SATN22]. GPU-Driven [KKRK22]. GPUs [DMB^{+}22, TPK^{+}18]. Gradient [CSX17, DBS17, SX19, WLS19]. Grained [XJ22, CMH^{+}17]. Graph [CKMP19, EC18, JJS17]. Graphs [DCKG19, HC19, LKE19, SZL19, YXL21]. Great [MSJ^{+}19, ZMW^{+}20]. Greedy [TTL^{+}21]. Ground [HHP20]. Ground-Truth [HHP20]. Growth [SVL20].

Guarantee [ZCW21]. Guarantees [FVB^{+}18, KCT18, LSNI21]. Guardrails [GSHB19].

Hadoop [GGYZ17]. Hardware

[GKK20, LYJ^{+}21]. Hardware-Assisted [GKK20]. Hardware-Aware [LYJ^{+}21]. Harvesting [AN20]. Hashing [SQ20]. Heart [KCMX20]. Heavy [GSHB19, SvKB^{+}20, Yin17, ZWT^{+}17, ZWT^{+}18, ZTS18]. Heavy-Tailed [SvKB^{+}20]. Heavy-Traffic [ZWT^{+}17, ZWT^{+}18, ZTS18]. Hegemony [ZHM^{+}21]. Heterogeneous [AG22, AAJV20, ZKAV20]. Hieroglyph [JJS17]. High [AMA18, HKY^{+}19, LKJK20, SX19]. High-Density [LKJK20]. High-Endurance [AMA18]. High-Performance [HKY^{+}19]. Homophilous [ZHM^{+}21]. hops [YXL21]. Host [WXW22]. Hound [ZL18]. Hybrid [ASME19]. Hyper [vdBBvL19]. Hyper-Scalable [vdBBvL19]. HyperBench [WZT19]. Hyperexponential [Van19].

I/O [AMA18, WXW22]. iBox [ATN^{+}22]. Identify [ZHZ^{+}21]. If [PSC19]. Imbalance [PJDQ20, ZWT^{+}18]. Impact [PST17, YHG^{+}17]. Implications [AGGS20]. Impression [KMM21]. Improving [AAJV20, GYSHB21, LGC^{+}18]. In-depth [CMH^{+}17]. Increasingly [ASV^{+}21]. Index [AdV18, LYL^{+}22]. Induced [LKS^{+}17]. Industrial [WLdA^{+}17]. Inequalities [ABG^{+}20]. Inference [Cas17, SMF^{+}22]. Inferring [BSA^{+}19, SK19]. Influence [NNVD17]. Information [BZ18, LCS^{+}21]. Infrastructures [SAM^{+}22]. Input [FKP^{+}21, GTL^{+}17, YNJS21]. Input-Dynamic [FKP^{+}21]. Input-Queued [GTL^{+}17]. Insensitivity [KV21]. Inserted [SDD^{+}17]. Instances [WUNK17]. Instrumentation [SDD^{+}17]. Intel [NBB^{+}19, OKB^{+}18]. Interacting [DE20]. Interactions [GLH^{+}19]. Interactive [AI19]. Internet [MSJ^{+}19, PJDQ20, SK19, ZMW^{+}20]. Intersection [ZZCC17]. Interval [IBL^{+}19]. Inventory [LYP^{+}19, YHS^{+}20]. Inverse [FS20]. Investigation [LML^{+}20, QWQ17]. Issues [TL20].

Job [SvKB^{+}20, Van19]. Jobs [WXHB21, WW20]. Join [HSV19, Zub20]. Join-Below-Threshold [HSV19]. JSQ [vdBBvL19].

Keeping [GSHB19]. Kernel [KCMX20]. Key [HKY^{+}19, SQ20]. Key-value [SQ20]. Kingman [CP18]. Knapsack [HCZ^{+}21, SZL^{+}20, TTL^{+}21, YZH^{+}21]. Know [FS20, IBL^{+}19, NBB^{+}19].

Labeling [DCKG19]. Lancet [CJC^{+}19]. Large [HV21, LML^{+}20, MDBvL17, SSM^{+}17]. Large-Scale [HV21, LML^{+}20, MDBvL17]. Last [IBL^{+}19]. Latency [LKS^{+}17, PJDQ20]. Law [YXL21]. layer [OKB^{+}18]. Learning [BSA^{+}21, CES19, CSX17, HC19, MPIZ17, RS22, SGS19, SBY19, SZL19, SX19, TP18, WCX^{+}18, WYN18, XLY^{+}22, YDH^{+}18, ZL18, Zho22]. Least [HV18]. Leave [ALBS^{+}20]. Length [KCT18, LM18]. Length-Bounded [KCT18]. Lessons [GYG^{+}18]. level [KB20]. Lifetime [CAJK17, LGC^{+}18]. Light [Yin17]. Lightweight [FVB^{+}18]. Like [IYRR18]. Limitation [ZWT^{+}18]. Limits [BBS20, FM19, WLS19]. Linear [FS20, LYQ^{+}22, LGK22, YHS^{+}20, Zho22]. Linux [QWQ17]. Live [SLWS20]. Load [ALR17, CNW^{+}17, GSHB19, HBV19, HV21, HSV19, MCS^{+}19, MBvL18, PJDQ20, PSC19, TS19, WZS20, YLX17, ZWT^{+}17, ZWT^{+}18, ZTS18]. Load-Balanced [PJDQ20, YLX17].

Loaded [HV18, HV21]. Loading [LYL^{+}22]. Loads [YLX17]. Local [SZL19]. Locality [TPK^{+}18, TKK21, WZS20]. Locally [JJS17]. Locally-Sufficient [JJS17]. Location [PTD20, SMF^{+}22, YWB19]. Location-Based [YWB19]. Logarithmic [LVW19]. Longevity [CAJK17]. Look [ASV^{+}21, SLJ19]. Look-Ahead [SLJ19]. Lookups [SQ20]. Loss [CPR19, LGC^{+}18]. Losses [LGW20]. Loud [IYRR18]. Low [AGS17, GSHB19, LYE19, LKJK20, TP18, WS17b, ZWT^{+}17]. Low-Complexity [AGS17, WS17b, ZWT^{+}17]. Low-Delay [WS17b]. Low-Wear [LKJK20]. LRU [QTES19, TQJS18]. LTE [PLX22, TLL^{+}18, VPK18]. LTE-Aware [VPK18]. LTERadar [VPK18]. Ludo [SQ20].

M [SGHB20, SHBSW20]. M/G/ [SGHB20]. M/G/1 [SHBSW20]. Machine [BSA^{+}21, CSX17]. Magma [HHP20]. Management [LNY21, QTES19, SIS17, YWH17, YHS^{+}20]. Many [BZ18]. Many-Sources [BZ18]. March [CCR22]. Marketplaces [ABG^{+}19]. Markov [CT19, DV18, WYN18, WKWY19]. Markov-Modulated [DV18]. Markovian [LKE19]. Match [ABG^{+}20, TKK21]. Matching [YXL21]. Matrix [AASS18, GFL^{+}22, MCS^{+}19]. Max [JIAM19]. Max-Min [JIAM19]. Maximization [TTL^{+}21]. Maximizing [DAR^{+}21]. Mean [AG22, Gas17, GV17, HV21, HSV19, RBT21, SSM^{+}17, Van19, Yin17]. Mean-Field [Gas17, SSM^{+}17, RBT21]. Measurement [BRC21, ZTY^{+}21, ZZCC17, ZZM^{+}19]. Measurements [JHWC21]. Measuring [PPL^{+}21, PTD20]. Mechanism [TSLG^{+}20]. Mechanisms [CEF^{+}20, CYG^{+}17, LKS^{+}17]. Meets [WCX^{+}18]. Membership [PTD20]. Memories [ASME19]. Memory [CAJK17, GFL^{+}22, LGC^{+}18, RKK22, SZL19, ZKAV20]. Merge [JJS17]. Merging [GKK20]. Message [WA18]. Metamorphic [XLY^{+}22]. Method [ALR17, Yin17]. Methodology [JKR18]. Methods [Cas17, DBS17]. Metric [SBY19]. Metrics [NT19]. Middle [ZYH^{+}19]. Millimeter [RS22]. Millimeter-Wave [RS22]. Min [JIAM19]. Mini [ZTY^{+}21]. Mini-Apps [ZTY^{+}21]. Minimal [IMMP20]. Minimization [LGK22]. Minimizing [LM18, QTES19]. Mirror [WYN20]. Mis [HWW^{+}20]. Misinformation [WKWY19]. Misses [QTES19]. Mix [TKK21]. Mixed [AI19]. Mobile [BZY^{+}22, BRC21, CNW^{+}17, TLL^{+}18]. Mobility [PLX22, SSB18]. Mobility-on-Demand [SSB18]. Model [AASS18, CMH^{+}17, KV21, LGH^{+}17, PST17, SSM^{+}17]. Modeling [RVC^{+}19, WLdA^{+}17]. Models [BSA^{+}19, GYG^{+}18, LM18, Van19]. Modem [CMH^{+}17]. Modern [CYG^{+}17, LKS^{+}17]. Modified [TTL^{+}21]. Modulated [DV18]. Monitoring [IBL^{+}19]. Monotone [TTL^{+}21]. Monte [LKE19]. MPX [OKB^{+}18]. mRSC [AMSS19]. Multi [AMSS19, IYRR18, LVW19, QTES19, SGS19, TS19, WXHB21, YPA^{+}18]. Multi-Agent [LVW19]. Multi-armed [YPA^{+}18]. Multi-dimensional [AMSS19]. Multi-flow [QTES19]. Multi-Server [TS19, WXHB21]. Multi-tenant [IYRR18]. Multicut [KCT18]. Multidimensional [YZH^{+}21]. Multilayer [JNTG18]. Multipartite [NVGT20]. Multiple [SZL^{+}20, WS17a]. Multiplication [GFL^{+}22, MCS^{+}19]. My [SIR20, BRC21].

N1 [CCR22]. N3 [CCR21]. Named [GGYZ17]. NAND [LKJK20, LGC^{+}18]. Near [MCS^{+}19, SHBSW20, TKZ^{+}18]. Near-Optimal [SHBSW20]. Near-Perfect [MCS^{+}19]. Nearly [SGHB20]. Necessary [ZTS18]. Network [ATN^{+}22, BBS20, BNS17, BSA^{+}21, CJC^{+}19, FWB17, FM19, IBL^{+}19, KCT18, LM18, NGB^{+}21, NT19, SAM^{+}22, SQ20, STGM21, VFV17, WS17b, WCX^{+}18, YHG^{+}17, ZZCC17, ZZM^{+}19]. Network-Caching [BBS20]. Networking [FVB^{+}18, GGYZ17, WXW22, ZGN^{+}19]. Networks [ALR17, AdV18, BZ18, FKP^{+}21, JNTG18, LSNI21, NNVD17, RGBY^{+}19, TLL^{+}18, TWFO20, TL20, WMX17, XJ22, ZLM^{+}18, ZHM^{+}21]. Neural [LYJ^{+}21, NGB^{+}21, WCX^{+}18]. NextG [XJ22]. NG [XJ22]. NG-Scope [XJ22]. no [ALBS^{+}20]. Node [PST17]. Noise [SIR20]. Noisy [HC19]. Non [DMB^{+}22, LKE19, LGW20, LGK22, PG17, WS17a, YDH^{+}18]. Non-additive [WS17a]. Non-Convex [YDH^{+}18, LGW20]. Non-Markovian [LKE19]. Non-Preemptive [PG17]. Non-stationary [LGK22]. Non-Uniform [DMB^{+}22]. Nonideal [TL20]. Nonlinear [PSLW22]. Nonlinearity [EC18]. Novel [LKJK20]. Nudge [GYSHB21]. NURA [DMB^{+}22].

O [AMA18, WXW22]. Obfuscating [ALBS^{+}20]. Object [BBHB18]. Observations [YPA^{+}18]. ODE [Van19]. ODE-based [Van19]. Offline [LNY21]. One [LYJ^{+}21, NKY^{+}18, SHBSW18]. Online [ABG^{+}19, AN20, BBS20, CYC^{+}19, LNY21, LCS^{+}21, LSNI21, LYP^{+}19, LGW20, MPIZ17, PSLW22, SLJ19, SZL^{+}20, TSLG^{+}20, TMM20, WYN18, WYN20, YWH17, YDH^{+}18, YHS^{+}20, YZH^{+}21, ZLW17]. Operating [KCMX20]. Operation [CYG^{+}17]. Optical [ZCW21]. Optimal [BBHB18, DV18, MDBvL17, MBvL18, SvKB^{+}20, SGHB20, SHBSW20, TMM20, WMX17, WKWY19, WZS20, YWH17, YDH^{+}18, ZLW17, ZKAV20, ZWT^{+}17, CPF20, Zub20]. Optimality [BDHB17, ZWT^{+}18, ZTS18]. Optimally [MPIZ17]. Optimization [AdV18, AN20, CYC^{+}19, DBS17, LYP^{+}19, LGW20, LVW19, PSLW22, SLJ19, WS17b, YHS^{+}20]. Optimizing [AI19]. Option [SZL19]. Order [CPF20]. Order-optimal [CPF20]. Outbound [MLHC20]. Outward [NNVD17]. Overcoming [ZWT^{+}18]. Overlapping [TMM20].

Packet [NPAP19]. Page [CNW^{+}17, LYL^{+}22, PSC19]. PageRank [VS19]. Paradigm [QTES19]. Parallel [BDHB17, LGH^{+}17, WW20]. Parallelism [TPK^{+}18]. Partial [CKMP19, Zub20]. Partition [LGH^{+}17]. Party [CEF^{+}20, PBLC21]. Patching [WLdA^{+}17]. Path [ATN^{+}22, STGM21]. Paths [PJDQ20]. Payment [TWFO20]. PDE [ALR17]. Peering [WM20]. Per-Flow [JIAM19]. Perfect [MCS^{+}19]. Performance [AAJV20, BCM17, Cas17, HKY^{+}19, HBV19, IKS21, NBB^{+}19, PBLC21, RVC^{+}19, SATN22, WXW22, ZCW21, ZGN^{+}19, ZMW^{+}20]. Persistent [ZZCC17]. Personalized [VS19]. Perspective [GZB^{+}21]. Physical [ZCW21]. Picocells [RS22]. Placement [KKRK22, PST17, ZKAV20]. Platforms [AMA18, BZY^{+}22]. Points [VPK18]. Policies [HBV19, SHBSW18, SvKB^{+}20, Zub20]. Policy [SGHB20]. Polynomial [CCG21]. Polynomial-Time [CCG21]. POMACS [CCR21, CCR22]. Ponzi [CLS^{+}21]. Pooling [TQJS18]. Pools [BCM17]. Portfolio [SIS17]. Portfolio-driven [SIS17]. Posted [ZLW17]. PoW [DAR^{+}21]. Power [CMH^{+}17, EC18, FWB17, GYG^{+}18, GZB^{+}21, HV18, HV21, IYRR18, SCYO21, YXL21]. Power-Law [YXL21]. Power-of-d-choices [HV18]. Practical [BBHB18, BSA^{+}19, SQ20]. Practicality [KCMX20]. Practices [JHWC21, WLdA^{+}17]. Predict [ABG^{+}20]. Predictable [RS22]. Prediction [NGB^{+}21, STGM21, ZLM^{+}18]. Predictions [CYC^{+}19, LYQ^{+}22, LGW20, LYE19]. Predictive [YHG^{+}17]. PredictRoute [STGM21]. Preemptive [PG17]. PreFix [ZLM^{+}18]. Price [KMM21, PPL^{+}21, SSB18]. Prices [ZLW17]. Pricing [SCYO21, WMX17]. Primal [WYN20]. Primal-Dual [WYN20]. Privacy [ALBS^{+}20, PPL^{+}21, PTD20, TWFO20]. Privacy-Utility [TWFO20]. Private [SAM^{+}22, WM20, Zho22]. Proactive [LYE19]. Problem [FS20, KCT18, SZL^{+}20]. Problems [YZH^{+}21]. Process [LGC^{+}18]. Processes [CES19, WYN18]. Processing [GFL^{+}22, JJS17]. Processing-In-Memory [GFL^{+}22]. Processor [KV21]. PROFET [RVC^{+}19]. Profile [ASME19]. Profile-Driven [ASME19]. Profiling [PSC19]. Program [SDD^{+}17]. Prophet [ABG^{+}20]. Proportional [FZCL18, GTL^{+}17]. Proportionally [TP18]. Providers [CEF^{+}20, WM20, YWB19]. Provisioning [CYC^{+}19]. Proxy [LYJ^{+}21]. Pruned [CJC^{+}19]. Public [WUNK17]. Pull [ZTS18]. Pull-based [ZTS18].

QoE [AdV18]. QoS [NT19, YWH17].

Quadratic [LYQ^{+}22]. Quality

[BNS17, BSA^{+}19]. Quantifying [TPK^{+}18]. Quantitative [JKR18]. Queries [IBL^{+}19]. Query [SATN22]. Queue [BZ18, CPR19, GTL^{+}17, LM18, ZWT^{+}18]. Queue-Proportional [GTL^{+}17]. Queued [GTL^{+}17]. Queueing [WXHB21, WW20]. Queues [CPR19]. Quickest [WKWY19]. QuickStop [WKWY19].

Radiometric [ALBS^{+}20]. Random [DE20, TS19, Zub20]. Randomized [ALR17, YWH17, YLX17]. Rankings [MPIZ17]. Rate [DBS17]. Rateless [MCS^{+}19]. Rates [JIAM19]. Rationing [ASME19]. Re [HNS19]. Re- [HNS19]. Reaching [KCT18]. Real [GFL^{+}22, KMM21, PHSB21]. Real-time [KMM21, PHSB21]. Receipts [NPAP19]. Reconfigurable [YCX^{+}19]. Recovery [CKMP19]. Recursive [BCM17]. Recycling [RKK22]. Redesigning [VFV17]. Reduced [CYG^{+}17]. Reduced-Voltage [CYG^{+}17]. Reduction [LKS^{+}17]. Redundancy [AAJV20, RB20, Zub20]. Refined [AG22, GV17]. Refinery [BSA^{+}21]. Refining [RBT21]. Regimes [Yin17]. Register [ZZCC17]. Regret [BBS20, CPF20, LM18, LGK22, TP18]. Regrets [CYC^{+}19]. Regulators [TL20]. Reinforcement [SBY19, Zho22]. Reliable [LKJK20]. Renewal [CES19]. Rényi [CKMP19, DCKG19]. Reorganizing [TKK21]. Replication [SSM^{+}17]. Representation [BSA^{+}21]. Reproducing [ZHM^{+}21]. Resilience [CJC^{+}19, KCT18, NT19, YNJS21]. Resilient [YHG^{+}17]. Resistances [YWB19]. resistant [RKTV22]. Resource [BCM17, CYC^{+}19, DMB^{+}22, HSV19, RKTV22, SIS17, TQJS18, TSLG^{+}20, ZLW17]. Resources [WS17a]. Response [GSHB19, SvKB^{+}20]. Restless [DV18]. Result [VS19]. Results [GGYZ17]. Retail [AS17]. Retention [LGC^{+}18]. Retroactive [NPAP19]. Revisiting [TTL^{+}21]. Richer [NGB^{+}21]. Road [GSHB19]. Robust [AMSS19]. Robustness [LYQ^{+}22]. Root [LML^{+}20]. Router [SDA18]. Routing [PHSB21, PST17, SK19, TWFO20, YCX^{+}19]. Rows [GKK20]. Ruin [CEF^{+}20]. Rules [ZHZ^{+}21].

SADPonzi [CLS^{+}21]. Safe [YLX17]. Same [GKK20]. Sampling [GTL^{+}17, NPAP19, TS19]. Scalability [GSWV20]. Scalable [vdBBvL19]. Scale [HV21, KCT18, LML^{+}20, MDBvL17, QWQ17, NNVD17, ZL18]. Scaling [BZ18]. Scam [XWG^{+}21]. Scenario [ZHZ^{+}21].

Scenario-dependent [ZHZ^{+}21]. Scheduling [BDHB17, GTL^{+}17, PG17, SHBSW18, SHBSW20]. Scheme [AMA18]. Schemes [CLS^{+}21, ZWT^{+}17]. Scope [XJ22]. Search [LYJ^{+}21]. Second [KMM21]. Securing [SX19]. Security [ALBS^{+}20, WS17a, ZYH^{+}19]. Segmentation [ABG^{+}19]. Segmentation-Thickness [ABG^{+}19]. Selection [HV18]. Semi [CCG21]. Semi-Bandits [CCG21]. Sensitive [TL20]. Separation [TQJS18]. Series [AASS18, PTD20]. Server [HV18, TS19, WXHB21]. Servers [HSV19, SIS17]. Service [LML^{+}20, MDBvL17, YWB19]. Services [LYE19, SSB18]. Set [KCMX20, NKY^{+}18]. Sets [CJC^{+}19]. Settings [CSX17]. SGD [CMM22]. SGX [NBB^{+}19]. Sharing [FZCL18, HSV19, KV21]. Should [NBB^{+}19]. Side [IYRR18]. Sided [WMX17]. Signal [SIR20]. Simple [CEF^{+}20, LGH^{+}17, SHBSW20]. Simulating [RVC^{+}19]. Simulation [ATN^{+}22, JKR18]. Size [NNVD17, NKY^{+}18]. Sizes [BBHB18, SvKB^{+}20, Van19]. Sizing [YNJS21]. Sketch [ZZM^{+}19]. Slowdowns [Zub20]. Smart [BRC21, CLS^{+}21, DAR^{+}21, PHSB21]. Smoothness [CPF20]. SOAP [SHBSW18]. Social [SGS19]. Software [ZHZ^{+}21]. Solving [FS20]. Some [IYRR18]. Sources [BZ18]. Space [RKK22]. Spaces [SBY19]. Sparse [GFL^{+}22, vdBBvL19]. SparseP [GFL^{+}22]. Spectral [CT19]. Speed [LYL^{+}22]. Speeding [YNJS21]. Spread [ZZCC17]. SSD [LNY21]. Stability [GSWV20, RB20]. Stack [OKB^{+}18, QWQ17]. Staleness [AEKB20]. State [BZ18, JIAM19, RBT21]. Stationary [CMM22, LGK22]. Statistical [CSX17, SX19]. Statistically [CCG21]. Stay [ALBS^{+}20]. Stealthy [SIR20]. Stein [Yin17]. Stepsize [CMM22]. Stochastic [JNTG18, SSM^{+}17, WYN20]. Stochastically [GYSHB21]. Stopping [WKWY19]. Storage [BRC21, CAJK17, LKJK20, SSM^{+}17, ZKAV20]. Store [HKY^{+}19]. Straggler [ZL18]. Strategies [PST17, TMM20]. Streaming [BZY^{+}22, BSA^{+}19, KLCS19, NGB^{+}21]. Structural [VS19]. Structure [RGBY^{+}19]. Structures [HKY^{+}19]. Study [FWB17, GYG^{+}18, GLH^{+}19, KKRK22, PLX22, SATN22, WYN18, ZTY^{+}21]. style [AN20]. Submodular [HCZ^{+}21, TTL^{+}21]. Substation [FWB17]. Sufficient [JJS17, ZTS18]. SUGAR [YNJS21]. Suite [WZT19]. Sum [DBS17]. Summarization [HCZ^{+}21]. Summer [IBL^{+}19]. Supermarket [KV21]. Supply [ABG^{+}20]. Supporting [DMB^{+}22, TLL^{+}18]. Surface [SLWS20]. Switch [NVGT20, ZLM^{+}18]. Switched [ZCW21]. Switches [GTL^{+}17]. Switching [PSLW22, SIR20, YLX17]. Sync [JJS17]. Synchronization [TL20]. Synthesis [SDA18]. Synthetic [AMSS19]. System [KCMX20, OKB^{+}18, RVC^{+}19]. Systematic [ZHZ^{+}21]. Systems [AG22, Cas17, GSWV20, HV21, KV21, LVW19, LGK22, MDBvL17, NBB^{+}19, PHSB21, RB20, SQ20, SSM^{+}17, TS19, WLdA^{+}17, ZKAV20, ZWT^{+}18, ZTS18, Zub20].

Tailed [SvKB^{+}20]. Tails [SvKB^{+}20]. Targeting [TMM20]. Tasks [TKK21]. TCP [MSJ^{+}19, QWQ17]. Techniques [NKY^{+}18]. TeksDB [HKY^{+}19]. Telemetry [XJ22]. Telling [GYG^{+}18]. tenant [IYRR18]. Testing [XLY^{+}22]. Theoretic [FZCL18]. Theory [KMM21, ZWT^{+}17]. Thermal [KKRK22]. Thermal-Aware [KKRK22]. Thickness [ABG^{+}19]. Third [CEF^{+}20, PBLC21]. Third-Party [CEF^{+}20, PBLC21]. Threading [KB20]. Threats [ZYH^{+}19]. Threshold [HSV19]. Throughput [GZB^{+}21]. Time [AASS18, CCG21, HV21, PSC19, PTD20, SvKB^{+}20, TL20, WYN18, KMM21, PHSB21]. Time-Sensitive [TL20]. Time-Series [PTD20]. Times [GSHB19]. Timing [IYRR18]. Tokens [XWG^{+}21]. Tolerant [SDA18]. Tolerating [LGC^{+}18]. Tomographic [PST17]. Too [KB20]. Toolkit [STGM21]. Topologies [MBvL18]. Topology [GZB^{+}21, WCX^{+}18, ZCW21]. Trace [ALBS^{+}20]. Traceroute [MLHC20]. Traces [AGGS20]. Tracking [GWX^{+}20]. Trade [XWG^{+}21]. Tradeoff [ABG^{+}19]. Tradeoffs [TWFO20]. Traffic [ASV^{+}21, AGGS20, BSA^{+}19, BSA^{+}21, GSHB19, NPAP19, WCX^{+}18, Yin17, ZWT^{+}17, ZZM^{+}19, ZWT^{+}18, ZTS18]. Traffic-driven [WCX^{+}18]. Transient [AI19, SIS17]. Transit [PHSB21]. Transnational [ZMW^{+}20]. Trick [XWG^{+}21]. Truncation [RBT21]. Truth [HHP20]. TTL [KLCS19]. TTL-based [KLCS19]. Tuxedo [DAR^{+}21]. Two [CP18, WMX17]. Two-Sided [WMX17]. Type [CMM22].

Uncertain [ABG^{+}20, LYE19].

Uncertainty [HC18]. Understanding [CYG^{+}17, HWW^{+}20, JHWC21, SVL20, SLWS20, WXW22, ZCW21, ZGN^{+}19]. Unified [TSLG^{+}20]. Uniform [DMB^{+}22]. Unimodal [CPF20]. Uniswap [XWG^{+}21]. Untrusted [LYQ^{+}22]. upon [GYSHB21]. Usefulness [SVL20]. User [KB20, PPL^{+}21]. User-level [KB20]. Using [IYRR18, PSC19, SZL19, WUNK17, YWB19, IBL^{+}19]. Utilities [WS17a]. Utility [TWFO20].

V5 [CCR21]. V6 [CCR22]. Value

[HKY^{+}19, SLJ19, SQ20]. Values [Gas17]. Variable [BBHB18]. Variation [LKS^{+}17, LGC^{+}18]. Vector [DE20, GFL^{+}22, MCS^{+}19]. Vehicle [SZL^{+}20, YWB19, ZHZ^{+}21]. via [ALBS^{+}20, AASS18, CKMP19, DE20, Gas17, JJS17, NT19]. Video [BSA^{+}19, NGB^{+}21].

View [PJDQ20]. Violations [ZHZ^{+}21]. Virtualization [WZT19, ZYH^{+}19]. Virtualized [AMA18, NBB^{+}19, NKY^{+}18]. Vision [RS22]. VM [PG17]. VMs [AI19]. Voltage [CYG^{+}17]. Volume [FM19]. Volume-based [FM19]. VR [TLL^{+}18]. vrfinder [MLHC20]. Vulnerability [QWQ17].

Waiting [HV21]. Walk [TS19]. Walks [DE20]. Wave [RS22]. Weakly [WYN18]. Wear [LKJK20, ZGN^{+}19]. Weaving [HKY^{+}19]. Web [ASV^{+}21, LYL^{+}22, PSC19]. Webcams [SLWS20]. Wechat [ZTY^{+}21]. Weighted [IMMP20]. What-If [PSC19]. Where [BRC21]. Whittle [AdV18]. Who [SVL20]. Wi [VPK18]. Wi-Fi [VPK18]. WiFi [ALBS^{+}20]. Wild [SLWS20]. Wireless [AdV18, BZ18]. Without [JIAM19, RVC^{+}19, BZ18, CPF20, PPL^{+}21]. Working [NKY^{+}18]. Workload [GLH^{+}19, HBV19, HNS19]. Workload-DRAM [GLH^{+}19]. Workloads [AI19]. Works [AG22]. Write [ASME19]. Write-Rationing [ASME19].

Xatu [NGB^{+}21].

YourAdvalue [PPL^{+}21].

Zero [WXHB21, WW20].

References

Anton:2020:IPH

[AAJV20] Elene Anton, Urtzi Ayesta, Matthieu Jonckheere, and Ina Maria Verloop. Improving the performance of heterogeneous data centers through redundancy. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (3):48:1–48:29, November 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3428333.

Agarwal:2018:MAT

[AASS18] Anish Agarwal, Muhammad Jehangir Amjad, Devavrat Shah, and Dennis Shen. Model agnostic time series analysis via matrix estimation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (3):40:1–40:39, December 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3287319.

Alijani:2019:STT

[ABG^{+}19] Reza Alijani, Siddhartha Banerjee, Sreenivas Gollapudi, Kostas Kollias, and Kamesh Munagala. The segmentation-thickness tradeoff in online marketplaces. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (1):18:1–18:26, March 2019. CO-DEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311089.

Alijani:2020:PMP

[ABG^{+}20] Reza Alijani, Siddhartha Banerjee, Sreenivas Gollapudi, Kamesh Munagala, and Kangning Wang. Predict and match: Prophet inequalities with uncertain supply. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (1):04:1–04:23, May 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3379470.

Anand:2018:WIB

[AdV18] Arjun Anand and Gustavo de Veciana. A Whittle’s index based approach for QoE optimization in wireless networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1): 15:1–15:39, April 2018. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179418.

Aral:2020:SCE

[AEKB20] Atakan Aral, Melike Erol-Kantarci, and Ivona Brandić. Staleness control for edge data analytics. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 4(2):38:1–38:24, June 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3392156.

Allmeier:2022:MFR

[AG22] Sebastian Allmeier and Nicolas Gast. Mean field and refined mean field approximations for heterogeneous systems: It works! Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(1):13:1–13:43, March 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508033.

Avin:2020:CTT

[AGGS20] Chen Avin, Manya Ghobadi, Chen Griner, and Stefan Schmid. On the complexity of traffic traces and implications. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (1):20:1–20:29, May 2020. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3379486.

Avrachenkov:2017:LCA

[AGS17] Konstantin Avrachenkov, Jasper Goseling, and Berksan Serbetci. A low-complexity approach to distributed cooperative caching with geographic constraints. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (1):27:1–27:??, June 2017. CO-DEN ???? ISSN 2476-

                                        1249. URL http://dl.acm.org/
                                        citation.cfm?id=3084465.

Ambati:2019:OCE

[AI19] Pradeep Ambati and David Irwin. Optimizing the cost of executing mixed interactive and batch workloads on transient VMs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):28:1–28:24, June 2019. CO-DEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3341617.3326143.

Abanto-Leon:2020:SCL

[ALBS^{+}20] Luis Fernando Abanto-Leon, Andreas Bäuml, Gek Hong (Allyson)

Sim, Matthias Hollick, and Arash Asadi. Stay connected, leave no trace: Enhancing security and privacy in WiFi via obfuscating radiometric fingerprints. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (3):44:1–44:31, November 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3428329.

Aghajani:2017:PMA

[ALR17] Reza Aghajani, Xingjie Li, and Kavita Ramanan. The PDE method for the analysis of randomized load balancing networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):38:1–38:28, December 2017. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3154497.

Ahmadian:2018:ECH

[AMA18] Saba Ahmadian, Onur Mutlu, and Hossein Asadi. ECI-Cache: a high-endurance and cost-efficient I/O caching scheme for virtualized platforms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1):9:1–9:34, April 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179412.

Amjad:2019:MMD

[AMSS19] Muhammad Amjad, Vishal Misra, Devavrat Shah, and Dennis Shen. mRSC: Multidimensional robust synthetic control. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):37:1–37:27, June 2019. CO-DEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3341617.3326152.

Asgari:2020:BSO

[AN20] Kamiar Asgari and Michael J. Neely. Bregman-style online convex optimization with energyharvesting constraints. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (3):52:1–52:25, November 2020. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3428337.

Amjad:2017:CDE

[AS17] Muhammad J. Amjad and Devavrat Shah. Censored demand estimation in retail. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (2):31:1–31:28, December 2017. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3154489.

Akram:2019:CGP

[ASME19] Shoaib Akram, Jennifer Sartor, Kathryn McKinley, and Lieven Eeckhout. Crystal gazer: Profiledriven write-rationing garbage collection for hybrid memories. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (1):9:1–9:27, March 2019. CO-DEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311080.

Akbari:2021:LBC

[ASV^{+}21] Iman Akbari, Mohammad A.

Salahuddin, Leni Ven, Noura Limam, Raouf Boutaba, Bertrand Mathieu, Stephanie Moteau, and Stephane Tuffin. A look behind the curtain: Traffic classification in an increasingly encrypted Web. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(1):04:1–04:26, February 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3447382.

Ashok:2022:DDN

[ATN^{+}22] Sachin Ashok, Shubham Tiwari, Nagarajan Natarajan, Venkata N. Padmanabhan, and Sundararajan Sellamanickam. Data-driven network path simulation with iBox. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1): 6:1–6:26, March 2022. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3508026.

Berger:2018:PBO

[BBHB18] Daniel S. Berger, Nathan Beckmann, and Mor Harchol-Balter. Practical bounds on optimal caching with variable object sizes. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(2):32:1–32:38, June 2018. CO-DEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3224427.

Bhattacharjee:2020:FLR

[BBS20] Rajarshi Bhattacharjee, Subhankar Banerjee, and Abhishek Sinha. Fundamental limits on the regret of online network-caching. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (2):25:1–25:31, June 2020. CO-DEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3392143.

Bonald:2017:PBF

[BCM17] Thomas Bonald, Céline Comte, and Fabien Mathieu. Performance of balanced fairness in resource pools: a recursive approach. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):41:1–41:25, December 2017. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3154500.

Bonald:2019:E

[BD19] Thomas Bonald and Nick Duffield. Editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):20:1– 20:2, June 2019. CODEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3341617.3326134.

Berg:2017:TOP

[BDHB17] Benjamin Berg, Jan-Pieter Dorsman, and Mor Harchol-Balter. Towards optimality in parallel scheduling. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):40:1– 40:30, December 2017. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3154499.

Bondorf:2017:QCD

[BNS17] Steffen Bondorf, Paul Nikolaus, and Jens B. Schmitt. Quality and cost of deterministic network calculus: Design and evaluation of an accurate and fast analysis. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 1(1):16:1–16:??, June 2017. CODEN ???? ISSN 2476-

1249. URL http://dl.acm.org/
citation.cfm?id=3084453.

Bijlani:2021:WDM

[BRC21] Ashish Bijlani, Umakishore Ramachandran, and Roy Campbell. Where did my 256 GB go? A measurement analysis of storage consumption on smart mobile devices. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 5(2):28:1–28:28, June 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3460095.

Bronzino:2019:ISV

[BSA^{+}19] Francesco Bronzino, Paul Schmitt, Sara Ayoubi, Guilherme Martins, Renata Teixeira, and Nick Feamster. Inferring streaming video quality from encrypted traffic: Practical models and deployment experience. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(3):56:1– 56:25, December 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3366704.

Bronzino:2021:TRC

[BSA^{+}21] Francesco Bronzino, Paul Schmitt, Sara Ayoubi, Hyojoon Kim, Renata Teixeira, and Nick Feamster. Traffic refinery: Cost-aware data representation for machine learning on network traffic. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):40:1–40:24, December 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491052.

Borst:2018:DSM

[BZ18] Sem Borst and Martin Zubeldia. Delay scaling in manysources wireless networks without queue state information. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (2):34:1–34:45, June 2018. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3224429.

Bhuyan:2022:EEC

[BZY^{+}22] Sandeepa Bhuyan, Shulin Zhao, Ziyu Ying, Mahmut T. Kandemir, and Chita R. Das. Endto-end characterization of game streaming applications on mobile platforms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(1):10:1–10:25, March 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508030.

Choi:2017:EDL

[CAJK17] Wonil Choi, Mohammad Arjomand, Myoungsoo Jung, and Mahmut Kandemir. Exploiting data longevity for enhancing the lifetime of flash-based storage class memory. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 1(1):21:1–21:??, June 2017. CODEN ???? ISSN 2476-

1249. URL http://dl.acm.org/
citation.cfm?id=3084458.

Casale:2017:API

[Cas17]  Giuliano Casale.  Accelerat-

ing performance inference over closed systems by asymptotic methods. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 1(1):8:1–8:??, June 2017. CODEN ???? ISSN 2476-

1249. URL http://dl.acm.org/
citation.cfm?id=3084445.

Cuvelier:2021:SEP

[CCG21] Thibaut Cuvelier, Richard Combes, and Eric Gourdin. Statistically efficient, polynomial-time algorithms for combinatorial semibandits. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(1):09:1–09:31, February 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3447387.

Carlsson:2021:PVN

[CCR21] Niklas Carlsson, Edith Cohen, and Philippe Robert. POMACS V5, N3, December 2021 editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(3):29:1, December 2021. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491041.

Carlsson:2022:PVN

[CCR22] Niklas Carlsson, Edith Cohen, and Philippe Robert. PO-MACS V6, N1, March 2022 editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(1):1:1, March 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508021.

Cai:2020:TPD

[CEF^{+}20] Yang Cai, Federico Echenique, Hu Fu, Katrina Ligett, Adam Wierman, and Juba Ziani. Third-party data providers ruin simple mechanisms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1):

12:1–12:31, May 2020. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3379478.

Cayci:2019:LCR

[CES19] Semih Cayci, Atilla Eryilmaz, and R. Srikant. Learning to control renewal processes with bandit feedback. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):43:1– 43:32, June 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3341617.3326158.

Chaintreau:2017:E

[CGZ17] Augustin Chaintreau, Leana Golubchik, and Zhi-Li Zhang. Editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 1(1):1:1–1:??, June 2017. CODEN ???? ISSN 2476-

1249. URL http://dl.acm.org/
citation.cfm?id=3105875.

Chang:2019:LBN

[CJC^{+}19] Yiyang Chang, Chuan Jiang, Ashish Chandra, Sanjay Rao, and Mohit Tawarmalani. Lancet: Better network resilience by designing for pruned failure sets. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):49:1–49:26, December 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366697.

Cullina:2019:PRE

[CKMP19] Daniel Cullina, Negar Kiyavash, Prateek Mittal, and H. Vincent Poor. Partial recovery of Erdős–Rényi graph alignment via k-core alignment. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):54:1–54:21, December 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366702.

Chen:2021:SDC

[CLS^{+}21] Weimin Chen, Xinran Li, Yuting Sui, Ningyu He, Haoyu Wang, Lei Wu, and Xiapu Luo. SADPonzi: Detecting and characterizing Ponzi schemes in Ethereum smart contracts. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (2):26:1–26:30, June 2021. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3460093.

Chen:2017:FGE

[CMH^{+}17] Xiaomeng Chen, Jiayi Meng, Y. Charlie Hu, Maruti Gupta, Ralph Hasholzner, Venkatesan Nallampatti Ekambaram, Ashish Singh, and Srikathyayani Srikanteswara. A fine-grained event-based modem power model for enabling in-depth modem energy drain analysis. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):45:1– 45:28, December 2017. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3154504.

Chen:2022:SBC

[CMM22] Zaiwei Chen, Shancong Mou, and Siva Theja Maguluri. Stationary behavior of constant stepsize SGD type algorithms: an asymptotic characterization. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1): 19:1–19:24, March 2022. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508039.

Cao:2017:DEC

[CNW^{+}17] Yi Cao, Javad Nejati, Muhammad Wajahat, Aruna Balasubramanian, and Anshul Gandhi. Deconstructing the energy consumption of the mobile page load. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):6:1–6:??, June 2017. CO-DEN ???? ISSN 2476-

1249. URL http://dl.acm.org/
citation.cfm?id=3084443.

Ciucu:2018:TEK

[CP18] Florin Ciucu and Felix Poloczek.

Two extensions of Kingman’s GI/G/1 bound. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(3):43:1– 43:33, December 2018. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3287322.

Combes:2020:UBC

[CPF20] Richard Combes, Alexandre Proutière, and Alexandre Fauquette. Unimodal bandits with continuous arms: Order-optimal regret without smoothness. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1): 14:1–14:28, May 2020. CODEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3379480.

Ciucu:2019:QLD

[CPR19] Florin Ciucu, Felix Poloczek, and Amr Rizk. Queue and loss distributions in finite-buffer queues. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (2):31:1–31:29, June 2019. CO-DEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3341617.3326146.

Chen:2017:DSM

[CSX17] Yudong Chen, Lili Su, and Jiaming Xu. Distributed statistical machine learning in adversarial settings: Byzantine gradient descent. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):44:1–44:25, December 2017. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3154503.

Combes:2019:CEE

[CT19] Richard Combes and Mikael Touati. Computationally efficient estimation of the spectral gap of a Markov chain. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1): 7:1–7:21, March 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311078.

Comden:2019:OOC

[CYC^{+}19] Joshua Comden, Sijie Yao, Niangjun Chen, Haipeng Xing, and Zhenhua Liu. Online optimization in cloud resource provisioning: Predictions, regrets, and algorithms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1): 16:1–16:30, March 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311087.

Chang:2017:URV

[CYG^{+}17] Kevin K. Chang, A. Giray Yaălikçi, Saugata Ghose, Aditya Agrawal, Niladrish Chatterjee, Abhijith Kashyap, Donghyuk Lee, Mike O’Connor, Hasan Hassan, and Onur Mutlu. Understanding reduced-voltage operation in modern DRAM devices: Experimental characterization, analysis, and mechanisms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):10:1– 10:??, June 2017. CODEN ???? ISSN 2476-1249. URL http://

dl.acm.org/citation.cfm?id=
3084447.

Das:2021:TMS

[DAR^{+}21] Sourav Das, Nitin Awathare, Ling Ren, Vinay J. Ribeiro, and Umesh Bellur. Tuxedo: Maximizing smart contract computation in PoW blockchains. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):41:1–41:30, December 2021. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3491053.

Doan:2017:CRD

[DBS17] Thinh T. Doan, Carolyn L. Beck, and R. Srikant. On the convergence rate of distributed gradient methods for finite-sum optimization under communication delays. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):37:1–37:27, December 2017. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3154496.

Dai:2019:ACL

[DCKG19] Osman Emre Dai, Daniel Cullina, Negar Kiyavash, and Matthias Grossglauser. Analysis of a canonical labeling algorithm for the alignment of correlated Erdős–Rényi graphs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2): 36:1–36:25, June 2019. CODEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3341617.3326151.

Doshi:2020:FVA

[DE20] Vishwaraj Doshi and Do Young Eun. Fiedler vector approximation via interacting random walks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1):01:1–01:28, May 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3379502.

Darabi:2022:NFS

[DMB^{+}22] Sina Darabi, Negin Mahani, Hazhir Baxishi, Ehsan Yousefzadeh-Asl-Miandoab, Mohammad Sadrosadati, and Hamid Sarbazi-Azad. NURA: a framework for supporting non-uniform resource accesses in GPUs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1): 16:1–16:27, March 2022. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3508036.

Duran:2018:AOC

[DV18] Santiago Duran and Ina Maria Verloop. Asymptotic optimal control of Markov-modulated restless bandits. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1):7:1–7:25, April 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179410.

Eliav:2018:BGD

[EC18] Buchnik Eliav and Edith Cohen. Bootstrapped graph diffusions: Exposing the power of nonlinearity. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 2(1):10:1–10:19, April 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179413.

Foerster:2021:IDD

[FKP^{+}21] Klaus-Tycho Foerster, Janne H. Korhonen, Ami Paz, Joel Rybicki, and Stefan Schmid. Inputdynamic distributed algorithms for communication networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (1):06:1–06:33, February 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3447384.

Fu:2019:FLV

[FM19] Xinzhe Fu and Eytan Modiano. Fundamental limits of volumebased network DoS attacks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):50:1–50:36, December 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366698.

Fleder:2019:FAD

[FS19] Michael Fleder and Devavrat Shah. Forecasting with alternative data. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(3):46:1– 46:29, December 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3366694.

Fleder:2020:KWY

[FS20] Michael Fleder and Devavrat Shah. I know what you bought at Chipotle for $9.81 by solving a linear inverse problem. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (3):47:1–47:17, November 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3428332.

Fanti:2018:DLC

[FVB^{+}18] Giulia Fanti, Shaileshh Bojja Venkatakrishnan, Surya Bakshi, Bradley Denby, Shruti Bhargava, Andrew Miller, and Pramod Viswanath. Dandelion++: Lightweight cryptocurrency networking with formal anonymity guarantees. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(2): 29:1–29:35, June 2018. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3224424.

Formby:2017:CSP

[FWB17] David Formby, Anwar Walid, and Raheem Beyah. A case study in power substation network dynamics. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 1(1):19:1–19:??, June 2017. CODEN ???? ISSN 2476-

                                        1249. URL http://dl.acm.org/
                                        citation.cfm?id=3084456.

Freeman:2018:DPS

[FZCL18] Rupert Freeman, Seyed Majid Zahedi, Vincent Conitzer, and Benjamin C. Lee. Dynamic proportional sharing: a gametheoretic approach. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1):3:1–3:36, April 2018. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3179406.

Gast:2017:EVE

[Gas17] Nicolas Gast. Expected values estimated via mean-field approximation are 1/N-accurate. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (1):17:1–17:??, June 2017. CO-DEN ???? ISSN 2476-

                                        1249. URL http://dl.acm.org/
                                        citation.cfm?id=3084454.

Giannoula:2022:STE

[GFL^{+}22] Christina Giannoula, Ivan Fernandez, Juan Gómez Luna, Nectarios Koziris, Georgios Goumas, and Onur Mutlu. SparseP: Towards efficient sparse matrix vector multiplication on real processing-in-memory architectures. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(1):21:1–21:49, March 2022. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3508041.

Gibbens:2017:HND

[GGYZ17] Mathias Gibbens, Chris Gniady, Lei Ye, and Beichuan Zhang. Hadoop on named data networking: Experience and results. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (1):2:1–2:??, June 2017. CO-DEN ???? ISSN 2476-

1249. URL http://dl.acm.org/
citation.cfm?id=3084439.

Ghahani:2020:DCH

[GKK20] Seyed Armin Vakil Ghahani, Mahmut Taylan Kandemir, and Jagadish B. Kotra. DSM: a case for hardware-assisted merging of DRAM rows with same content. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 4(2):33:1–33:26, June 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3392151.

Gandhi:2021:E

[GKW21] Anshul Gandhi, Negar Kiyavash, and Jia Wang. Editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (2):13:1, June 2021. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3466793.

Ghose:2019:DCW

[GLH^{+}19] Saugata Ghose, Tianshi Li, Nastaran Hajinazar, Damla Senol Cali, and Onur Mutlu. Demystifying complex workload-DRAM interactions: an experimental study. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(3):60:1–60:50, December 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366708.

Grosof:2019:LBG

[GSHB19] Isaac Grosof, Ziv Scully, and Mor Harchol-Balter. Load balancing guardrails: Keeping your heavy traffic on the road to low response times. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2): 42:1–42:31, June 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3341617.3326157.

Gopalan:2020:SSB

[GSWV20] Aditya Gopalan, Abishek Sankararaman, Anwar Walid, and Sriram Vishwanath. Stability and scalability of blockchain systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2): 35:1–35:35, June 2020. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3392153.

Gong:2017:QPS

[GTL^{+}17] Long Gong, Paul Tune, Liang Liu, Sen Yang, and Jun (Jim) Xu. Queue-proportional sampling: A better approach to crossbar scheduling for inputqueued switches. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):3:1–3:??, June 2017. CODEN ???? ISSN 2476-

1249. URL http://dl.acm.org/
citation.cfm?id=3084440.

Gast:2017:RMF

[GV17] Nicolas Gast and Benny Van Houdt. A refined mean field approximation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):33:1– 33:28, December 2017. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3154491.

Gao:2020:TCC

[GWX^{+}20] Bingyu Gao, Haoyu Wang, Pengcheng Xia, Siwei Wu, Yajin Zhou, Xiapu Luo, and Gareth Tyson. Tracking counterfeit cryptocurrency end-to-end. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (3):50:1–50:28, November 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3428335.

Ghose:2018:WYD

[GYG^{+}18] Saugata Ghose, Abdullah Giray Yaglikçi, Raghav Gupta, Donghyuk Lee, Kais Kudrolli, William X. Liu, Hasan Hassan, Kevin K. Chang, Niladrish Chatterjee, Aditya Agrawal, Mike O’Connor, and Onur Mutlu. What your DRAM power models are not telling you: Lessons from a detailed experimental study. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (3):38:1–38:41, December 2018. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3224419.

Grosof:2021:NSI

[GYSHB21] Isaac Grosof, Kunhe Yang, Ziv Scully, and Mor Harchol-Balter. Nudge: Stochastically improving upon FCFS. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 5(2):21:1–21:29, June 2021. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3460088.

Griner:2021:CPC

[GZB^{+}21] Chen Griner, Johannes Zerwas, Andreas Blenk, Manya Ghobadi, Stefan Schmid, and Chen Avin. Cerberus: The power of choices in datacenter topology design — a throughput perspective. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):38:1–38:33, December 2021. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3491050.

Hellemans:2019:PAW

[HBV19] Tim Hellemans, Tejas Bodas, and Benny Van Houdt. Performance analysis of workload dependent load balancing policies. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (2):35:1–35:35, June 2019. CO-DEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3341617.3326150.

Hoffmann:2018:CUC

[HC18] Jessica Hoffmann and Constantine Caramanis. The cost of uncertainty in curing epidemics. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (2):31:1–31:33, June 2018. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3224426.

Hoffmann:2019:LGN

[HC19] Jessica Hoffmann and Constantine Caramanis. Learning graphs from noisy epidemic cascades. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (2):40:1–40:34, June 2019. CO-DEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3341617.3326155.

Han:2021:AAS

[HCZ^{+}21] Kai Han, Shuang Cui, Tianshuai Zhu, Enpei Zhang, Benwei Wu, Zhizhuo Yin, Tong Xu, Shaojie Tang, and He Huang. Approximation algorithms for submodular data summarization with a knapsack constraint. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (1):05:1–05:31, February 2021. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3447383.

Hazimeh:2020:MGT

[HHP20] Ahmad Hazimeh, Adrian Herrera, and Mathias Payer. Magma: a ground-truth fuzzing benchmark. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(3):49:1–49:29, November 2020. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3428334.

Han:2019:TWD

[HKY^{+}19] Youil Han, Bryan S. Kim, Jeseong Yeon, Sungjin Lee, and Eunji Lee. TeksDB: Weaving data structures for a highperformance key–value store. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (1):8:1–8:23, March 2019. CO-DEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3322205.3311079.

Henzinger:2019:EDW

[HNS19] Monika Henzinger, Stefan Neumann, and Stefan Schmid. Efficient distributed workload (re- )Embedding. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):13:1– 13:38, March 2019. CODEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3322205.3311084.

Horvath:2019:MFA

[HSV19] Illés Antal Horváth, Ziv Scully, and Benny Van Houdt. Mean field analysis of join-below-threshold load balancing for resource sharing servers. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):57:1–57:21, December 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366705.

Hellemans:2018:PDC

[HV18] Tim Hellemans and Benny Van Houdt. On the power-of-d-choices with least loaded server selection. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 2(2):27:1–27:22, June 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3224422.

Hellemans:2021:MWT

[HV21] Tim Hellemans and Benny Van Houdt. Mean waiting time in large-scale and critically loaded power of d load balancing systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2):19:1–19:34, June 2021. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3460086.

Huang:2020:UMB

[HWW^{+}20] Yuheng Huang, Haoyu Wang, Lei Wu, Gareth Tyson, Xiapu Luo, Run Zhang, Xuanzhe Liu, Gang Huang, and Xuxian Jiang. Understanding (mis)Behavior on the EOSIO blockchain. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2): 37:1–37:28, June 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3392155.

Ivkin:2019:KWY

[IBL^{+}19] Nikita Ivkin, Ran Ben Basat, Zaoxing Liu, Gil Einziger, Roy Friedman, and Vladimir Braverman. I know what you did last summer: Network monitoring using interval queries. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):61:1–61:28, December 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366709.

Iqbal:2021:DCG

[IKS21] Hassan Iqbal, Ayesha Khalid, and Muhammad Shahzad. Dissecting cloud gaming performance with DECAF. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):31:1–31:27, December 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491043.

Im:2020:DWF

[IMMP20] Sungjin Im, Benjamin Moseley, Kamesh Munagala, and Kirk Pruhs. Dynamic weighted fairness with minimal disruptions. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (1):19:1–19:18, May 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3379485.

Islam:2018:WSL

[IYRR18] Mohammad A. Islam, Luting Yang, Kiran Ranganath, and Shaolei Ren. Why some like it loud: Timing power attacks in multi-tenant data centers using an acoustic side channel. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (1):6:1–6:33, April 2018. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179409.

Jin:2021:UPG

[JHWC21] Lin Jin, Shuai Hao, Haining Wang, and Chase Cotton. Understanding the practices of global censorship through accurate, end-to-end measurements. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):43:1–43:25, December 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491055.

Jose:2019:DAC

[JIAM19] Lavanya Jose, Stephen Ibanez, Mohammad Alizadeh, and Nick McKeown. A distributed algorithm to calculate max-min fair rates without per-flow state. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2): 21:1–21:42, June 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3341617.3326135.

Ju:2017:HLS

[JJS17] Xiaoen Ju, Hani Jamjoom, and Kang G. Shin. Hieroglyph: Locally-sufficient graph processing via compute–sync–merge. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (1):9:1–9:??, June 2017. CODEN ???? ISSN 2476-1249. URL

http://dl.acm.org/citation.
cfm?id=3084446.

Jain:2018:QEC

[JKR18] Akshay Jain, Mahmoud Khairy, and Timothy G. Rogers. A quantitative evaluation of contemporary GPU simulation methodology. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(2):35:1–35:28, June 2018. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3224430.

Jiang:2018:CSM

[JNTG18] Bo Jiang, Philippe Nain, Don Towsley, and Saikat Guha. On a class of stochastic multilayer networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 2(1):18:1–18:25, April 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179421.

Karsten:2020:ULT

[KB20] Martin Karsten and Saman Barghi. User-level threading: Have your cake and eat it too. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (1):17:1–17:30, May 2020. CO-DEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3379483.

Kuo:2020:SCH

[KCMX20] Hsuan-Chi Kuo, Jianyan Chen, Sibin Mohan, and Tianyin Xu. Set the configuration for the heart of the OS: On the practicality of operating system kernel debloating. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 4(1):03:1–03:27, May 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3379469.

Kuhnle:2018:NRL

[KCT18] Alan Kuhnle, Victoria G. Crawford, and My T. Thai. Network resilience and the lengthbounded multicut problem: Reaching the dynamic billionscale with guarantees. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1): 4:1–4:26, April 2018. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3179407.

Khadirsharbiyani:2022:DCG

[KKRK22] Soheil Khadirsharbiyani, Jagadish Kotra, Karthik Rao, and Mahmut Kandemir. Data convection: a GPU-driven case study for thermal-aware data placement in 3D DRAMs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1): 7:1–7:25, March 2022. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3508027.

Karakoy:2019:AAA

[KKT^{+}19] Mustafa Karakoy, Orhan Kislal, Xulong Tang, Mahmut Taylan Kandemir, and Meenakshi Arunachalam. Architectureaware approximate computing. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (2):38:1–38:24, June 2019. CO-DEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3341617.3326153.

Kumar:2019:TBA

[KLCS19] Dhruv Kumar, Jian Li, Abhishek Chandra, and Ramesh Sitaraman. A TTL-based approach for data aggregation in geo-distributed streaming analytics. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):29:1–29:27, June 2019. CO-DEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3341617.3326144.

Kinnear:2021:RTB

[KMM21] Ryan J. Kinnear, Ravi R.

Mazumdar, and Peter Marbach. Real-time bidding for time constrained impression contracts in first and second price auctions — theory and algorithms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):37:1–37:37, December 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491049.

Kielanski:2021:AIS

[KV21] Grzegorz Kielanski and Benny Van Houdt. On the asymptotic insensitivity of the supermarket model in processor sharing systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 5(2):22:1–22:28, June 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3460089.

Li:2021:IAC

[LCS^{+}21] Tongxin Li, Yue Chen, Bo Sun, Adam Wierman, and Steven H. Low. Information aggregation for constrained online control. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (2):18:1–18:35, June 2021. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3460085.

Luo:2018:INF

[LGC^{+}18] Yixin Luo, Saugata Ghose, Yu Cai, Erich F. Haratsch, and Onur Mutlu. Improving 3D NAND flash memory lifetime by tolerating early retention loss and process variation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(3):37:1– 37:48, December 2018. CODEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3224432.

Li:2017:SYE

[LGH^{+}17] Lingda Li, Robel Geda, Ari B. Hayes, Yanhao Chen, Pranav Chaudhari, Eddy Z. Zhang, and Mario Szegedy. A simple yet effective balanced edge partition model for parallel computing. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (1):14:1–14:??, June 2017. CO-DEN ???? ISSN 2476-

                                        1249. URL http://dl.acm.org/
                                        citation.cfm?id=3084451.

Luo:2022:DRM

[LGK22] Yuwei Luo, Varun Gupta, and Mladen Kolar. Dynamic regret minimization for control of nonstationary linear dynamical systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1):9:1–9:72, March 2022. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3508029.

Lin:2020:OOP

[LGW20] Yiheng Lin, Gautam Goel, and Adam Wierman. Online optimization with predictions and non-convex losses. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1): 18:1–18:32, May 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3379484.

Lee:2019:NMM

[LKE19] Chul-Ho Lee, Min Kang, and Do Young Eun. Non-Markovian Monte Carlo on directed graphs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (1):15:1–15:31, March 2019. CO-DEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311086.

Liu:2020:CNA

[LKJK20] Chun-Yi Liu, Jagadish Kotra, Myoungsoo Jung, and Mahmut Taylan Kandemir. Centaur: a novel architecture for reliable, low-wear, high-density 3D NAND storage. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2): 28:1–28:25, June 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3392146.

Lee:2017:DIL

[LKS^{+}17] Donghyuk Lee, Samira Khan, Lavanya Subramanian, Saugata Ghose, Rachata Ausavarungnirun, Gennady Pekhimenko, Vivek Seshadri, and Onur Mutlu. Design-induced latency variation in modern DRAM chips: Characterization, analysis, and latency reduction mechanisms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):26:1– 26:??, June 2017. CODEN ???? ISSN 2476-1249. URL http://

dl.acm.org/citation.cfm?id=
3084464.

Liang:2018:MQL

[LM18] Qingkai Liang and Eytan Modiano. Minimizing queue length regret under adversarial network models. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 2(1):11:1–11:32, April 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179414.

Lin:2020:FDA

[LML^{+}20] Fred Lin, Keyur Muzumdar, Nikolay Pavlovich Laptev, Mihai-Valentin Curelea, Seunghak Lee, and Sriram Sankar. Fast dimensional analysis for root cause investigation in a largescale service environment. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2): 31:1–31:23, June 2020. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3392149.

Lange:2021:OOA

[LNY21] Tomer Lange, Joseph (Seffi)

Naor, and Gala Yadgar. Offline and online algorithms for SSD management. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):33:1–33:28, December 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491045.

Li:2021:OCN

[LSNI21] Yuanyuan Li, Tareq Si Salem, Giovanni Neglia, and Stratis Ioannidis. Online caching networks with adversarial guarantees. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(3):35:1–35:39, December 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491047.

London:2019:LCD

[LVW19] Palma London, Shai Vardi, and Adam Wierman. Logarithmic communication for distributed optimization in multi-agent systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(3):48:1–48:29, December 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366696.

Liu:2019:PCL

[LYE19] Ran Liu, Edmund Yeh, and Atilla Eryilmaz. Proactive caching for low access-delay services under uncertain predictions. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):2:1–2:46, March 2019. CO-DEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311073.

Lu:2021:OPD

[LYJ^{+}21] Bingqian Lu, Jianyi Yang, Weiwen Jiang, Yiyu Shi, and Shaolei Ren. One proxy device is enough for hardware-aware neural architecture search. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):34:1–34:34, December 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491046.

Liu:2022:FSI

[LYL^{+}22] Wei Liu, Xinlei Yang, Hao Lin, Zhenhua Li, and Feng Qian. Fusing speed index during Web page loading. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(1):23:1–23:23, March 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3511214.

Lin:2019:COO

[LYP^{+}19] Qiulin Lin, Hanling Yi, John Pang, Minghua Chen, Adam Wierman, Michael Honig, and Yuanzhang Xiao. Competitive online optimization under inventory constraints. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1): 10:1–10:28, March 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311081.

Li:2022:RCL

[LYQ^{+}22] Tongxin Li, Ruixiao Yang, Guannan Qu, Guanya Shi, Chenkai Yu, Adam Wierman, and Steven Low. Robustness and consistency in linear quadratic control with untrusted predictions. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1): 18:1–18:35, March 2022. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508038.

Mukherjee:2018:AOL

[MBvL18] Debankur Mukherjee, Sem C.

Borst, and Johan S. H. van Leeuwaarden. Asymptotically optimal load balancing topologies. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1):14:1–14:29, April 2018. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179417.

Mallick:2019:RCN

[MCS^{+}19] Ankur Mallick, Malhar Chaudhari, Utsav Sheth, Ganesh Palanikumar, and Gauri Joshi. Rateless codes for near-perfect load balancing in distributed matrix–vector multiplication.

Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):58:1–58:40, December 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366706.

Mukherjee:2017:OSE

[MDBvL17] Debankur Mukherjee, Souvik Dhara, Sem C. Borst, and Johan S. H. van Leeuwaarden. Optimal service elasticity in largescale distributed systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):25:1– 25:??, June 2017. CODEN ???? ISSN 2476-1249. URL http://

dl.acm.org/citation.cfm?id=
3084463.

Marder:2020:VFO

[MLHC20] Alexander Marder, Matthew Luckie, Bradley Huffaker, and KC Claffy. vrfinder: Finding outbound addresses in Traceroute. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2):40:1–40:28, June 2020. CO-DEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3392158.

Magureanu:2017:OLO

[MPIZ17] Stefan Magureanu, Alexandre Proutiere, Marcus Isaksson, and Boxun Zhang. Online learning of optimally diverse rankings. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (2):32:1–32:26, December 2017.

CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3154490.

Mishra:2019:GIT

[MSJ^{+}19] Ayush Mishra, Xiangpeng Sun, Atishya Jain, Sameer Pande, Raj Joshi, and Ben Leong. The Great Internet TCP Congestion Control Census. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):45:1–45:24, December 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366693.

Ngoc:2019:EYS

[NBB^{+}19] Tu Dinh Ngoc, Bao Bui, Stella Bitchebe, Alain Tchana, Valerio Schiavoni, Pascal Felber, and Daniel Hagimont. Everything you should know about Intel SGX performance on virtualized systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):5:1–5:21, March 2019. CO-DEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311076.

Nam:2021:XRN

[NGB^{+}21] Yun Seong Nam, Jianfei Gao, Chandan Bothra, Ehab Ghabashneh, Sanjay Rao, Bruno Ribeiro, Jibin Zhan, and Hui Zhang. Xatu: Richer neural network based prediction for video streaming. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(3):44:1– 44:26, December 2021. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3491056.

Nitu:2018:WSS

[NKY^{+}18] Vlad Nitu, Aram Kocharyan, Hannas Yaya, Alain Tchana, Daniel Hagimont, and Hrachya Astsatryan. Working set size estimation techniques in virtualized environments: One size does not fit all. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 2(1):19:1–19:22, April 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179422.

Nguyen:2017:OIC

[NNVD17] Hung T. Nguyen, Tri P. Nguyen, Tam N. Vu, and Thang N. Dinh. Outward influence and cascade size estimation in billion-scale networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 1(1):20:1–20:??, June 2017. CODEN ???? ISSN 2476-

1249. URL http://dl.acm.org/
citation.cfm?id=3084457.

Nikolopoulos:2019:RPS

[NPAP19] Pavlos Nikolopoulos, Christos Pappas, Katerina Argyraki, and Adrian Perrig. Retroactive packet sampling for traffic receipts. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):19:1– 19:39, March 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311090.

Nguyen:2019:NRA

[NT19] Lan N. Nguyen and My T. Thai.

Network resilience assessment via QoS degradation metrics: an algorithmic approach. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1): 1:1–1:32, March 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311072.

Nain:2020:AME

[NVGT20] Philippe Nain, Gayane Vardoyan, Saikat Guha, and Don Towsley. On the analysis of a multipartite entanglement distribution switch. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 4(2):23:1–23:39, June 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3392141.

Oleksenko:2018:IME

[OKB^{+}18] Oleksii Oleksenko, Dmitrii Kuvaiskii, Pramod Bhatotia, Pascal Felber, and Christof Fetzer. Intel MPX explained: a crosslayer analysis of the Intel MPX system stack. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 2(2):28:1–28:30, June 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3224423.

Pourghassemi:2021:ACP

[PBLC21] Behnam Pourghassemi, Jordan Bonecutter, Zhou Li, and Aparna Chandramowlishwaran. adPerf: Characterizing the performance of third-party ads. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(1): 03:1–03:26, February 2021. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3447381.

Psychas:2017:NPV

[PG17] Konstantinos Psychas and Javad Ghaderi. On non-preemptive VM scheduling in the cloud. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (2):35:1–35:29, December 2017. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3154493.

Perivier:2021:RTA

[PHSB21] Noémie Périvier, Chamsi Hssaine, Samitha Samaranayake, and Siddhartha Banerjee. Realtime approximate routing for smart transit systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2): 24:1–24:30, June 2021. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3460091.

Pi:2020:LIA

[PJDQ20] Yibo Pi, Sugih Jamin, Peter Danzig, and Feng Qian. Latency imbalance among Internet load-balanced paths: a cloudcentric view. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 4(2):32:1–32:29, June 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3392150.

Pan:2022:FLC

[PLX22] Yueyang Pan, Ruihan Li, and Chenren Xu. The first 5G-LTE comparative study in extreme mobility. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(1):20:1–20:22, March 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508040.

Pachilakis:2021:YMA

[PPL^{+}21] Michalis Pachilakis, Panagiotis Papadopoulos, Nikolaos Laoutaris, Evangelos P. Markatos, and Nicolas Kourtellis. YourAdvalue: Measuring advertising price dynamics without bankrupting user privacy. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):32:1–32:26, December 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491044.

Pourghassemi:2019:WIA

[PSC19] Behnam Pourghassemi, Ardalan Amiri Sani, and Aparna Chandramowlishwaran. What-if analysis of page load time in Web browsers using causal profiling. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2): 27:1–27:23, June 2019. CODEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3341617.3326142.

Pan:2022:OOF

[PSLW22] Weici Pan, Guanya Shi, Yiheng Lin, and Adam Wierman. Online optimization with feedback delay and nonlinear switching cost. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(1):17:1–17:34, March 2022. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3508037.

Pignolet:2017:TNP

[PST17] Yvonne-Anne Pignolet, Stefan Schmid, and Gilles Tredan. Tomographic node placement strategies and the impact of the routing model. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):42:1– 42:23, December 2017. CODEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3154501.

Pyrgelis:2020:MMP

[PTD20] Apostolos Pyrgelis, Carmela Troncoso, and Emiliano De Cristofaro. Measuring membership privacy on aggregate location time-series. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2): 36:1–36:28, June 2020. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3392154.

Quan:2019:NFM

[QTES19] Guocong Quan, Jian Tan, Atilla Eryilmaz, and Ness Shroff. A new flexible multi-flow LRU cache management paradigm for minimizing misses. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2): 39:1–39:30, June 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3341617.3326154.

Quach:2017:ILT

[QWQ17] Alan Quach, Zhongjie Wang, and Zhiyun Qian. Investigation of the 2016 Linux TCP stack vulnerability at scale. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):4:1–4:??, June 2017. CODEN ???? ISSN 2476-

1249. URL http://dl.acm.org/
citation.cfm?id=3084441.

Raaijmakers:2020:ASR

[RB20] Youri Raaijmakers and Sem Borst. Achievable stability in redundancy systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(3):46:1– 46:21, November 2020. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3428331.

Randone:2021:RMF

[RBT21] Francesca Randone, Luca Bortolussi, and Mirco Tribastone. Refining mean-field approximations by dynamic state truncation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2):25:1–25:30, June 2021. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3460092.

Ros-Giralt:2019:BSC

[RGBY^{+}19] Jordi Ros-Giralt, Atul Bohara, Sruthi Yellamraju, M. Harper Langston, Richard Lethin, Yuang Jiang, Leandros Tassiulas, Josie Li, Yuanlong Tan, and Malathi Veeraraghavan. On the bottleneck structure of congestioncontrolled networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(3):59:1– 59:31, December 2019. CODEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3366707.

Ryoo:2022:MSR

[RKK22] Jihyun Ryoo, Mahmut Taylan Kandemir, and Mustafa Karakoy. Memory space recycling. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(1):14:1–14:24, March 2022. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3508034.

Rana:2022:FAR

[RKTV22] Ranvir Rana, Sreeram Kannan, David Tse, and Pramod Viswanath. Free2Shard: Adversaryresistant distributed resource allocation for blockchains. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1): 11:1–11:38, March 2022. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3508031.

Regmi:2022:APM

[RS22] Hem Regmi and Sanjib Sur. Argus: Predictable millimeter-wave picocells with vision and learning augmentation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(1):2:1–2:26, March 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508022.

Radulovic:2019:PMS

[RVC^{+}19] Milan Radulovic, Rommel Sánchez Verdejo, Paul Carpenter, Petar Radojković, Bruce Jacob, and Eduard Ayguadé. PROFET: Modeling system performance and energy without simulating the CPU. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):34:1– 34:33, June 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3341617.3326149.

Salamatian:2022:CBA

[SAM^{+}22] Loqman Salamatian, Scott Anderson, Joshua Matthews, Paul Barford, Walter Willinger, and Mark Crovella. Curvature-based analysis of network connectivity in private backbone infrastructures. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 6(1):5:1–5:32, March 2022. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508025.

Suh:2022:CES

[SATN22] Young-Kyoon Suh, Junyoung An, Byungchul Tak, and Gap-Joo Na. A comprehensive empirical study of query performance across GPU DBMSes. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1): 4:1–4:29, March 2022. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3508024.

Sinclair:2019:ADE

[SBY19] Sean R. Sinclair, Siddhartha Banerjee, and Christina Lee Yu. Adaptive discretization for episodic reinforcement learning in metric spaces. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):55:1–55:44, December 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366703.

Shin:2021:PBP

[SCYO21] Suho Shin, Hoyong Choi, Yung Yi, and Jungseul Ok. Power of bonus in pricing for crowdsourcing. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(3):36:1–36:25, December 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491048.

Subramanian:2018:SFT

[SDA18] Kausik Subramanian, Loris D’Antoni, and Aditya Akella. Synthesis of fault-tolerant distributed router configurations. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (1):22:1–22:26, April 2018. CO-DEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3179425.

Schardl:2017:CFC

[SDD^{+}17] Tao B. Schardl, Tyler Denniston, Damon Doucet, Bradley C. Kuszmaul, I-Ting Angelina Lee, and Charles E. Leiserson. The CSI framework for compilerinserted program instrumentation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):43:1–43:25, December 2017. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3154502.

Scully:2020:GPN

[SGHB20] Ziv Scully, Isaac Grosof, and Mor Harchol-Balter. The Gittins policy is nearly optimal in the M/G/k under extremely general conditions. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (3):43:1–43:29, November 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3428328.

Sankararaman:2019:SLM

[SGS19] Abishek Sankararaman, Ayalvadi Ganesh, and Sanjay Shakkottai. Social learning in multi agent multi armed bandits. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(3):53:1–53:35, December 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366701.

Scully:2018:SOC

[SHBSW18] Ziv Scully, Mor Harchol-Balter, and Alan Scheller-Wolf. SOAP: One clean analysis of all agebased scheduling policies. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1): 16:1–16:30, April 2018. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179419.

Scully:2020:SNO

[SHBSW20] Ziv Scully, Mor Harchol-Balter, and Alan Scheller-Wolf. Simple near-optimal scheduling for the M/G/1. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 4(1):11:1–11:29, May 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3379477.

Shao:2020:YNM

[SIR20] Zhihui Shao, Mohammad A. Islam, and Shaolei Ren. Your noise, my signal: Exploiting switching noise for stealthy data exfiltration from desktop computers. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1):07:1–07:39, May 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3379473.

Sharma:2017:PDR

[SIS17] Prateek Sharma, David Irwin, and Prashant Shenoy. Portfoliodriven resource management for transient cloud servers. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (1):5:1–5:??, June 2017. CO-DEN ???? ISSN 2476-

1249. URL http://dl.acm.org/
citation.cfm?id=3084442.

Sermpezis:2019:ICI

[SK19] Pavlos Sermpezis and Vasileios Kotronis. Inferring catchment in Internet routing. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2): 30:1–30:31, June 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3341617.3326145.

Shi:2019:VLA

[SLJ19] Ming Shi, Xiaojun Lin, and Lei Jiao. On the value of look-ahead in competitive online convex optimization. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2): 22:1–22:42, June 2019. CODEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3341617.3326136.

Song:2020:UCS

[SLWS20] JinKe Song, Qiang Li, Haining Wang, and Limin Sun. Under the concealing surface: Detecting and understanding live webcams in the wild. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 4(1):05:1–05:25, May 2020. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3379471.

Silva:2022:AIB

[SMF^{+}22] Brivaldo A. Silva, Paulo Mol, Osvaldo Fonseca, Italo Cunha, Ronaldo A. Ferreira, and Ethan Katz-Bassett. Automatic inference of BGP location communities. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1):3:1–3:23, March 2022. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3508023.

Shi:2020:LHC

[SQ20] Shouqian Shi and Chen Qian. Ludo Hashing: Compact, fast, and dynamic key-value lookups for practical network systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (2):22:1–22:32, June 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3392140.

Sejourne:2018:PFM

[SSB18] Thibault Séjournè, Samitha Samaranayake, and Siddhartha Banerjee. The price of fragmentation in mobility-on-demand services. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(2):30:1–30:26, June 2018. CO-DEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3224425.

Sun:2017:ASM

[SSM^{+}17] Wen Sun, Veronique Simon, Sebastien Monnet, Philippe Robert, and Pierre Sens. Analysis of a stochastic model of replication in large distributed storage systems: A mean-field approach. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):24:1–24:??, June 2017. CO-DEN ???? ISSN 2476-1249. URL

http://dl.acm.org/citation.
cfm?id=3084462.

Singh:2021:PNP

[STGM21] Rachee Singh, David Tench, Phillipa Gill, and Andrew Mc-Gregor. PredictRoute: a network path prediction toolkit. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (2):23:1–23:24, June 2021. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3460090.

Scully:2020:CPO

[SvKB^{+}20] Ziv Scully, Lucas van Kreveld, Onno Boxma, Jan-Pieter Dorsman, and Adam Wierman. Characterizing policies with optimal response time tails under heavy-tailed job sizes. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2): 30:1–30:33, June 2020. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3392148.

Snyder:2020:WFF

[SVL20] Peter Snyder, Antoine Vastel, and Ben Livshits. Who filters the filters: Understanding the growth, usefulness and efficiency of crowdsourced ad blocking. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 4(2):26:1–26:24, June 2020. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3392144.

Su:2019:SDG

[SX19] Lili Su and Jiaming Xu. Securing distributed gradient descent in high dimensional statistical learning. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):12:1– 12:41, March 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311083.

Su:2019:CLB

[SZL19] Lili Su, Martin Zubeldia, and Nancy Lynch. Collaboratively learning the best option on graphs, using bounded local memory. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):11:1– 11:32, March 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311082.

Sun:2020:CAO

[SZL^{+}20] Bo Sun, Ali Zeynali, Tongxin Li, Mohammad Hajiesmaili, Adam Wierman, and Danny H. K. Tsang. Competitive algorithms for the online multiple knapsack problem with application to electric vehicle charging. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (3):51:1–51:32, November 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3428336.

Tang:2021:MMR

[TKK21] Xulong Tang, Mahmut Taylan Kandemir, and Mustafa Karakoy. Mix and match: Reorganizing tasks for enhancing data locality. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 5(2):20:1–20:24, June 2021. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3460087.

Tang:2018:CND

[TKZ^{+}18] Xulong Tang, Mahmut Taylan Kandemir, Hui Zhao, Myoungsoo Jung, and Mustafa Karakoy. Computing with near data. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (3):42:1–42:30, December 2018. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3287321.

Thomas:2020:TSI

[TL20] Ludovic Thomas and Jean-Yves Le Boudec. On time synchronization issues in time-sensitive networks with regulators and nonideal clocks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 4(2):27:1–27:41, June 2020. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3392145.

Tan:2018:SMV

[TLL^{+}18] Zhaowei Tan, Yuanjie Li, Qianru Li, Zhehui Zhang, Zhehan Li, and Songwu Lu. Supporting mobile VR in LTE networks: How close are we? Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1):8:1–8:31, April 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179411.

Tay:2020:E

[TM20] Y. C. Tay and Athina Markopoulou.

Editorial. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 4(2):21:1, June 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3392139.

Tillberg:2020:OBS

[TMM20] Erik Tillberg, Peter Marbach, and Ravi Mazumdar. Optimal bidding strategies for online ad auctions with overlapping targeting criteria. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 4(2):34:1–34:55, June 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3392152.

Talebi:2018:LPF

[TP18] Mohammad Sadegh Talebi and Alexandre Proutiere. Learning proportionally fair allocations with low regret. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(2): 36:1–36:31, June 2018. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3224431.

Tang:2018:QDL

[TPK^{+}18] Xulong Tang, Ashutosh Pattnaik, Onur Kayiran, Adwait Jog, Mahmut Taylan Kandemir, and Chita Das. Quantifying data locality in dynamic parallelism in GPUs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(3):39:1–39:24, December 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3287318.

Tan:2018:RPS

[TQJS18] Jian Tan, Guocong Quan, Kaiyi Ji, and Ness Shroff. On resource pooling and separation for LRU caching. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1):5:1–5:31, April 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179408.

Tang:2019:RWB

[TS19] Dengwang Tang and Vijay G. Subramanian. Random walk based sampling for load balancing in multi-server systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1): 14:1–14:44, March 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311085.

Tan:2020:MDO

[TSLG^{+}20] Xiaoqi Tan, Bo Sun, Alberto Leon-Garcia, Yuan Wu, and Danny H. K. Tsang. Mechanism design for online resource allocation: a unified approach. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (2):24:1–24:46, June 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3392142.

Tang:2021:RMG

[TTL^{+}21] Jing Tang, Xueyan Tang, Andrew Lim, Kai Han, Chongshou Li, and Junsong Yuan. Revisiting modified greedy algorithm for monotone submodular maximization with a knapsack constraint. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(1):08:1–08:22, February 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3447386.

Tang:2020:PUT

[TWFO20] Weizhao Tang, Weina Wang, Giulia Fanti, and Sewoong Oh. Privacy-utility tradeoffs in routing cryptocurrency over payment channel networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (2):29:1–29:39, June 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3392147.

VanHoudt:2019:GAO

[Van19] Benny Van Houdt. Global attraction of ODE-based mean field models with hyperexponential job sizes. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):23:1–23:23, June 2019. CO-DEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3341617.3326137.

vanderBoor:2019:HSJ

[vdBBvL19] Mark van der Boor, Sem Borst, and Johan van Leeuwaarden. Hyper-scalable JSQ with sparse feedback. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1):4:1– 4:37, March 2019. CODEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3322205.3311075.

Venkatakrishnan:2017:DRB

[VFV17] Shaileshh Bojja Venkatakrishnan, Giulia Fanti, and Pramod Viswanath. Dandelion: Redesigning the Bitcoin network for anonymity. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 1(1):22:1–22:??, June 2017. CODEN ???? ISSN 2476-

                                        1249. URL http://dl.acm.org/
                                        citation.cfm?id=3084459.

Vlachou:2018:LTL

[VPK18] Christina Vlachou, Ioannis Pefkianakis, and Kyu-Han Kim. LTERadar: Towards LTE-aware Wi-Fi access points. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(2): 33:1–33:31, June 2018. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3224428.

Vial:2019:SRP

[VS19] Daniel Vial and Vijay Subramanian. A structural result for personalized PageRank and its algorithmic consequences. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2): 25:1–25:88, June 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3341617.3326140.

Wierman:2018:ME

[WA18] Adam Wierman and Aditya Akella. Message from the Editors. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(2):23:1, June 2018. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3224418.

Wang:2018:NNM

[WCX^{+}18] Mowei Wang, Yong Cui, Shihan Xiao, Xin Wang, Dan Yang, Kai Chen, and Jun Zhu. Neural network meets DCN: Trafficdriven topology adaptation with deep learning. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 2(2):26:1–26:25, June 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3224421.

Wei:2019:QMO

[WKWY19] Honghao Wei, Xiaohan Kang, Weina Wang, and Lei Ying. QuickStop: a Markov optimal stopping approach for quickest misinformation detection. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2): 41:1–41:25, June 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3341617.3326156.

Wang:2017:CMP

[WLdA^{+}17] Brandon Wang, Xiaoye Li, Leandro P. de Aguiar, Daniel S. Menasche, and Zubair Shafiq. Characterizing and modeling patching practices of industrial control systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):18:1– 18:??, June 2017. CODEN ???? ISSN 2476-1249. URL http://

dl.acm.org/citation.cfm?id=
3084455.

Wang:2019:FLA

[WLS19] Sinong Wang, Jiashang Liu, and Ness Shroff. Fundamental limits of approximate gradient coding. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(3):52:1–52:22, December 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366700.

Wang:2020:PPA

[WM20] Xin Wang and Richard T. B. Ma. On private peering agreements between content and access providers: a contractual equilibrium analysis. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (3):41:1–41:32, November 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3428326.

Wang:2017:OTS

[WMX17] Xin Wang, Richard T. B. Ma, and Yinlong Xu. On optimal two-sided pricing of congested networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):7:1–7:??, June 2017. CODEN ???? ISSN 2476-

1249. URL http://dl.acm.org/
citation.cfm?id=3084444.

Wang:2017:SGN

[WS17a] Sinong Wang and Ness Shroff.

Security game with non-additive utilities and multiple attacker resources. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 1(1):13:1–13:??, June 2017. CODEN ???? ISSN 2476-

1249. URL http://dl.acm.org/
citation.cfm?id=3084450.

Wang:2017:TFC

[WS17b] Sinong Wang and Ness Shroff.

Towards fast-convergence, lowdelay and low-complexity network optimization. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (2):34:1–34:32, December 2017. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3154492.

Wang:2017:UBI

[WUNK17] Cheng Wang, Bhuvan Urgaonkar, Neda Nasiriani, and George Kesidis. Using burstable instances in the public cloud: Why, when and how? Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(1):11:1– 11:??, June 2017. CODEN ???? ISSN 2476-1249. URL http://

                                        dl.acm.org/citation.cfm?id=
                                        3084448.

Weng:2020:AZA

[WW20] Wentao Weng and Weina Wang. Achieving zero asymptotic queueing delay for parallel jobs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (3):42:1–42:36, November 2020. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3428327.

Wang:2021:ZQM

[WXHB21] Weina Wang, Qiaomin Xie, and Mor Harchol-Balter. Zero queueing for multi-server jobs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (1):07:1–07:25, February 2021. CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3447385.

Wang:2022:UDC

[WXW22] Minhu Wang, Mingwei Xu, and Jianping Wu. Understanding I/O direct cache access performance for end host networking. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6(1): 22:1–22:37, March 2022. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3508042.

Wei:2018:OLW

[WYN18] Xiaohan Wei, Hao Yu, and Michael J. Neely. Online learning in weakly coupled Markov decision processes: a convergence time study. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 2(1):12:1–12:38, April 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179415.

Wei:2020:OPD

[WYN20] Xiaohan Wei, Hao Yu, and Michael J. Neely. Online primal-dual mirror descent under stochastic constraints. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(2): 39:1–39:36, June 2020. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3392157.

Weng:2020:OLB

[WZS20] Wentao Weng, Xingyu Zhou, and R. Srikant. Optimal load balancing with locality constraints. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (3):45:1–45:37, November 2020.

CODEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3428330.

Wei:2019:HBS

[WZT19] Song Wei, Kun Zhang, and Bibo Tu. HyperBench: a benchmark suite for virtualization capabilities. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):24:1–24:22, June 2019. CO-DEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3341617.3326138.

Xie:2022:NSF

[XJ22] Yaxiong Xie and Kyle Jamieson. NG-Scope: Fine-grained telemetry for NextG cellular networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (1):12:1–12:26, March 2022. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3508032.

Xiao:2022:MTD

[XLY^{+}22] Dongwei Xiao, Zhibo LIU, Yuanyuan Yuan, Qi Pang, and Shuai Wang. Metamorphic testing of deep learning compilers. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (1):15:1–15:28, March 2022. CO-DEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3508035.

Xia:2021:TTD

[XWG^{+}21] Pengcheng Xia, Haoyu Wang, Bingyu Gao, Weihang Su, Zhou Yu, Xiapu Luo, Chao Zhang, Xusheng Xiao, and Guoai Xu. Trade or trick?: Detecting and characterizing scam tokens on Uniswap decentralized exchange. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (3):39:1–39:26, December 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491051.

Yang:2019:AER

[YCX^{+}19] Zhenjie Yang, Yong Cui, Shihan Xiao, Xin Wang, Minming Li, Chuming Li, and Yadong Liu. Achieving efficient routing in reconfigurable DCNs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):47:1–47:30, December 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366695.

Yang:2018:OAO

[YDH^{+}18] Lin Yang, Lei Deng, Mohammad H. Hajiesmaili, Cheng Tan, and Wing Shing Wong. An optimal algorithm for online non-convex learning. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(2): 25:1–25:25, June 2018. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3224420.

Yang:2017:PIA

[YHG^{+}17] Sen Yang, Yan He, Zihui Ge, Dongmei Wang, and Jun Xu. Predictive impact analysis for designing a resilient cellular backhaul network. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):30:1– 30:33, December 2017. CODEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3154488.

Yang:2020:OLO

[YHS^{+}20] Lin Yang, Mohammad H. Hajiesmaili, Ramesh Sitaraman, Adam Wierman, Enrique Mallada, and Wing S. Wong. Online linear optimization with inventory management constraints. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1): 16:1–16:29, May 2020. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3379482.

Ying:2017:SMM

[Yin17] Lei Ying. Stein’s method for mean field approximations in light and heavy traffic regimes. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1 (1):12:1–12:??, June 2017. CO-DEN ???? ISSN 2476-

                                        1249. URL http://dl.acm.org/
                                        citation.cfm?id=3084449.

Yang:2017:SRL

[YLX17] Sen Yang, Bill Lin, and Jun Xu.

Safe randomized load-balanced switching by diffusing extra loads. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):29:1–29:37, December 2017. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3154487.

Yang:2021:SSG

[YNJS21] Lishan Yang, Bin Nie, Adwait Jog, and Evgenia Smirni. SUGAR: Speeding up GPGPU application resilience estimation with input sizing. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(1): 01:1–01:29, February 2021. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3447375.

Yun:2018:MAB

[YPA^{+}18] Donggyu Yun, Alexandre Proutiere, Sumyeong Ahn, Jinwoo Shin, and Yung Yi. Multi-armed bandit with additional observations. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (1):13:1–13:22, April 2018. CO-DEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3179416.

Yu:2019:ALB

[YWB19] Haoran Yu, Ermin Wei, and Randall A. Berry. Analyzing location-based advertising for vehicle service providers using effective resistances. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1): 6:1–6:35, March 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311077.

Yang:2017:ORO

[YWH17] Lin Yang, Wing Shing Wong, and Mohammad H. Hajiesmaili. An optimal randomized online algorithm for QoS buffer management. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):36:1–36:26, December 2017. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3154494.

Yu:2021:PDH

[YXL21] Liren Yu, Jiaming Xu, and Xiaojun Lin. The power of D-hops in matching power-law graphs. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (2):27:1–27:43, June 2021. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3460094.

Yang:2021:CAO

[YZH^{+}21] Lin Yang, Ali Zeynali, Mohammad H. Hajiesmaili, Ramesh K. Sitaraman, and Don Towsley. Competitive algorithms for online multidimensional knapsack problems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(3):30:1– 30:30, December 2021. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3491042.

Zhao:2021:UPG

[ZCW21] Shizhen Zhao, Peirui Cao, and Xinbing Wang. Understanding the performance guarantee of physical topology design for optical circuit switched data centers. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(3):42:1–42:24, December 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3491054.

Zhu:2019:UNP

[ZGN^{+}19] Xiao Zhu, Yihua Ethan Guo, Ashkan Nikravesh, Feng Qian, and Z. Morley Mao. Understanding the networking performance of Wear OS. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(1): 3:1–3:25, March 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311074.

Zhang:2021:CHE

[ZHM^{+}21] Yiguang Zhang, Jessy Xinyi Han, Ilica Mahajan, Priyanjana Bengani, and Augustin Chaintreau. Chasm in hegemony: Explaining and reproducing disparities in homophilous networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5 (2):16:1–16:38, June 2021. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3460083.

Zhou:2022:DPR

[Zho22] Xingyu Zhou. Differentially private reinforcement learning with linear function approximation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 6 (1):8:1–8:27, March 2022. CO-DEN ???? ISSN 2476-

                                        1249. URL https://dl.acm.
                                        org/doi/10.1145/3508028.

Zhang:2021:SFI

[ZHZ^{+}21] Qingzhao Zhang, David Ke Hong, Ze Zhang, Qi Alfred Chen, Scott Mahlke, and Z. Morley Mao. A systematic framework to identify violations of scenariodependent driving rules in autonomous vehicle software. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(2): 15:1–15:25, June 2021. CODEN ???? ISSN 2476-1249. URL

                                        https://dl.acm.org/doi/10.
                                        1145/3460082.

Zhang:2020:ODP

[ZKAV20] Lei Zhang, Reza Karimi, Irfan Ahmad, and Ymir Vigfusson. Optimal data placement for heterogeneous cache, memory, and storage systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4 (1):06:1–06:27, May 2020. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3379472.

Zheng:2018:HCL

[ZL18] Pengfei Zheng and Benjamin C.

Lee. Hound: Causal learning for datacenter-scale straggler diagnosis. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 2(1):17:1–17:36, April 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179420.

Zhang:2018:PSF

[ZLM^{+}18] Shenglin Zhang, Ying Liu, Weibin Meng, Zhiling Luo, Jiahao Bu, Sen Yang, Peixian Liang, Dan Pei, Jun Xu, Yuzhi Zhang, Yu Chen, Hui Dong, Xianping Qu, and Lei Song. PreFix: Switch failure prediction in datacenter networks. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1):2:1–2:29, April 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179405.

Zhang:2017:OPP

[ZLW17] Zijun Zhang, Zongpeng Li, and Chuan Wu. Optimal posted prices for online cloud resource allocation. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 1(1):23:1–23:??, June 2017. CODEN ???? ISSN 2476-

1249. URL http://dl.acm.org/
citation.cfm?id=3084460.

Zarchy:2019:ACC

[ZMSS19] Doron Zarchy, Radhika Mittal, Michael Schapira, and Scott Shenker. Axiomatizing congestion control. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3(2):33:1– 33:33, June 2019. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3341617.3326148.

Zhu:2020:CTI

[ZMW^{+}20] Pengxiong Zhu, Keyu Man, Zhongjie Wang, Zhiyun Qian, Roya Ensafi, J. Alex Halderman, and Haixin Duan. Characterizing transnational Internet performance and the Great Bottleneck of China. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 4(1):13:1–13:23, May 2020. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3379479.

Zhou:2018:HTD

[ZTS18] Xingyu Zhou, Jian Tan, and Ness Shroff. Heavy-traffic delay optimality in pull-based load balancing systems: Necessary and sufficient conditions. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2 (3):41:1–41:33, December 2018. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3287323.

Zhang:2021:MSW

[ZTY^{+}21] Yue Zhang, Bayan Turkistani, Allen Yuqing Yang, Chaoshun Zuo, and Zhiqiang Lin. A measurement study of Wechat mini-apps. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 5(2):14:1–14:25, June 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3460081.

Zubeldia:2020:DOP

[Zub20] Martin Zubeldia. Delay-optimal policies in partial fork-join systems with redundancy and random slowdowns. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 4(1):02:1– 02:49, May 2020. CODEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3379468.

Zhou:2017:DLC

[ZWT^{+}17] Xingyu Zhou, Fei Wu, Jian Tan, Yin Sun, and Ness Shroff. Designing low-complexity heavytraffic delay-optimal load balancing schemes: Theory to algorithms. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 1(2):39:1–39:30, December 2017. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3154498.

Zhou:2018:DQI

[ZWT^{+}18] Xingyu Zhou, Fei Wu, Jian Tan, Kannan Srinivasan, and Ness Shroff. Degree of queue imbalance: Overcoming the limitation of heavy-traffic delay optimality in load balancing systems. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 2(1):21:1–21:41, April 2018. CO-DEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3179424.

Zhang:2019:AMD

[ZYH^{+}19] Lei Zhang, Zhemin Yang, Yuyu He, Mingqi Li, Sen Yang, Min Yang, Yuan Zhang, and Zhiyun Qian. App in the middle: Demystify application virtualization in Android and its security threats. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (1):17:1–17:24, March 2019. CO-DEN ???? ISSN 2476-1249. URL

https://dl.acm.org/doi/10.
1145/3322205.3311088.

Zhou:2017:PSM

[ZZCC17] You Zhou, Yian Zhou, Min Chen, and Shigang Chen. Persistent spread measurement for big network data based on register intersection. Proceedings of the ACM on Measurement and Analysis of Computing Systems (PO-MACS), 1(1):15:1–15:??, June 2017. CODEN ???? ISSN 2476-

1249. URL http://dl.acm.org/
citation.cfm?id=3084452.

Zhu:2021:FBG

[ZZLL21] Zhaowei Zhu, Jingxuan Zhu, Ji Liu, and Yang Liu. Federated Bandit: a gossiping approach. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 5(1):02:1–02:29, February 2021. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3447380.

Zhou:2019:GSF

[ZZM^{+}19] You Zhou, Youlin Zhang, Chaoyi Ma, Shigang Chen, and Olufemi O. Odegbile. Generalized sketch families for network traffic measurement. Proceedings of the ACM on Measurement and Analysis of Computing Systems (POMACS), 3 (3):51:1–51:34, December 2019. CODEN ???? ISSN 2476-

1249. URL https://dl.acm.
org/doi/10.1145/3366699.